### v1

# The equilibrium properties of obvious strategy profiles

Abstract
This paper studies the equilibrium properties of the “obvious strategy profile” in large finite-player games. Each player in such a strategy profile simply adopts a randomized strategy as she would have used in a symmetric equilibrium of an idealized large game. We show that, under a continuity assumption, (i) obvious strategy profiles constitute a convergent sequence of approximate symmetric equilibria as the number of players tends to infinity, and (ii) realizations of such strategy profiles also form a convergent sequence of (pure strategy) approximate equilibria with probability approaching one. Our findings offer a solution that is easily implemented without coordination issues and is asymptotically optimal for players in large finite games. Additionally, we present a convergence result for approximate symmetric equilibria.
JEL classification: C60; C62; C72
Keywords: Large games, large finite games, auxiliary mapping, obvious strategy profile, symmetric equilibrium, approximate equilibrium.
1 Introduction
Games with a continuum of agents (referred to as large games) have been extensively studied in the literature. Typically, such games model strategic interactions involving a large but finite number of participants. By assuming a continuum of players where each player’s unilateral deviation does not affect the aggregate action distribution, various desirable properties, such as the existence of pure strategy Nash equilibrium, have been established in these idealized large games. Despite significant progress in recent decades, the predominant focus in existing literature has been on the existence and convergence of pure strategy Nash equilibria in large games. However, an important question remains unanswered: do players truly benefit from such an idealized process? Specifically, while players may acknowledge that the Nash equilibrium of the idealized large game predicts a desirable social choice outcome, how do they behave to implement such an outcome in their real-life (finite) games? Furthermore, given an implementation strategy profile, do players have incentives to adhere to this designated strategy profile?
The first question seems to have a straightforward answer: each player could simply adopt the strategy they would select in an idealized large game. For instance, if a Nash equilibrium in the large game requires every player to choose a common strategy, then players in the finite real game can follow that equilibrium strategy. While this approach works in some situations, it often fails due to coordination issues arising from multiple individual-optimal choices.
To illustrate this, consider a simple example involving many players in a holiday village who must decide between going to the beach (action $B$) or staying at the hotel (action $H$). For each player, the payoff for staying at the hotel is always 00. Conversely, if a player chooses to go to the beach, her payoff depends on the proportion of players who also choose the beach. Specifically, her payoff is positive if only a few players go to the beach; however, as more players join her there, her payoff decreases and eventually turns negative once the number of beachgoers becomes sufficiently large. Formally, let $I$ denote the set of players and $A=\{B,H\}$ represent the action set available to them. Let $\tau(B)$ (resp. $\tau(H)$) denote the proportion of players selecting action $B$ (resp. $H$). Players in this game share a common payoff function given by:
$u(H,\tau)=0,\,\,\,\,\,u(B,\tau)=\tfrac{1}{2}-\tau(B),$
where $\tau=(\tau(B),\tau(H))$ is an action distribution.
This is a large finite-player game, and to simplify the analysis, players assume they are participating in a game with a continuum of players. In this imagined large game, there exists a pure strategy Nash equilibrium where each action is chosen by exactly half of the players, as each player in this scenario is indifferent between the actions $H$ and $B$ [^1] It can be verified that this is the unique pure strategy Nash equilibrium. Therefore, the action distribution $\tau^{*}=(\tfrac{1}{2},\tfrac{1}{2})$ represents the desirable social choice outcome predicted by the large game model. However, issues arise when considering the implementation of this outcome in a real finite game. Although players are aware of the equilibrium action distribution, each player has multiple optimal choices available given this distribution, [^2] Given the equilibrium action distribution $\tau^{*}=(\tfrac{1}{2},\tfrac{1}{2})$, it follows that $u(H,\tau^{*})=u(B,\tau^{*})$, indicating that both $B$ and $H$ are best responses for each player. but they cannot arbitrarily select one action, as doing so would lead to a non-equilibrium outcome. Thus, if players in the real finite game aim to use pure strategies to achieve the outcome $(\tfrac{1}{2},\tfrac{1}{2})$, they must either communicate with one another before the game begins or rely on a social planner to divide them into two groups and designate an action for each group. However, both of these approaches are unrealistic in scenarios involving many players.
We address this coordination issue in two steps. Firstly, we observe that players in a game are distinguished by their characteristics, typically represented by their payoff functions. For example, drivers in a large traffic network are differentiated by the types of vehicles they drive, each associated with distinct travel times (i.e., different payoff functions). Similarly, buyers in a large online market can be categorized into groups based on their preferences or payoffs for different commodities. Thus, the coordination issue can be resolved if players make decisions based on their characteristics, allowing those with the same characteristic to make the same choice. Such an idea is captured by the notion of symmetry. Therefore, the first step of our approach is to symmetrize the equilibrium outcome. Specifically, given an equilibrium action distribution $\tau^{*}$, we can transform it into a symmetric Nash equilibrium $f^{*}$ using an auxiliary mapping. This mapping ensures that (i) players sharing the same characteristic (i.e., the same payoff function) choose the same strategy in $f^{*}$, and (ii) the action distribution of $f^{*}$ equals $\tau^{*}$.
Secondly, we introduce the concept of an “obvious strategy profile” in real finite-player games. In this strategy profile, players make decisions based on their characteristics, where each player selects a strategy associated with her characteristic in the symmetric equilibrium $f^{*}$ of the imagined large game. For illustration, consider the Beach vs. Hotel game introduced earlier. Since all players in this game share a common payoff function, they possess the same characteristic and consequently choose the same strategy in the equilibrium $f^{*}$. Specifically, all players in $f^{*}$ could adopt a simple randomized strategy $\mu^{*}$, choosing each action with equal probability. The law of large numbers then ensures that the realized action distribution from this symmetric strategy profile matches the equilibrium action distribution, with each action chosen by half of the players. Finally, in real-life finite scenarios, the obvious strategy profile assigns $\mu^{*}$ to each player.
Given the implementation via symmetrization and obvious strategy profile described above, a natural question arises: do players in the real-life finite game have an incentive to follow the obvious strategy profile? This question pertains to the equilibrium property of such a strategy profile. In our motivating example above, the answer is clear: obvious strategy profile forms an approximate Nash equilibrium: suppose that the number of players is $n$. Based on some combinatorial calculations, we can determine that each player’s expected payoff when choosing action $B$ (resp. action $H$) is $-\frac{1}{2n}$ (resp. 00). [^3] Since there are $n$ players in this game and each of them takes a proportion of $\frac{1}{n}$, the expected payoff for each player is $\sum_{k=0}^{n-1}(\tfrac{1}{2}-\tfrac{k+1}{n})\binom{n-1}{k}(\tfrac{1}{2})^{n-1}=\tfrac{1}{2}-\frac{n+1}{2n}=-\tfrac{1}{2n}$ by choosing action $B$ , and 00 by choosing action $H$. Therefore, the obvious strategy profile adopted by all the players is a $\tfrac{1}{4n}$-Nash equilibrium in the $n$-player game. [^4] In an $\varepsilon$-Nash equilibrium, a large portion of players (more than $1-\varepsilon$) choose strategies that are within $\varepsilon$ of their optimal payoffs; see Definition 4 below. Moreover, as $n$ tends to infinity, we observe a convergent sequence of approximate equilibria, indicating that adopting the obvious strategy profile is asymptotically optimal for each player.
However, the asymptotic optimality of obvious strategy profile does not hold in general situations. Example 1 in Section 3 presents a large game $G$ with a Nash equilibrium $g$; however, the obvious strategy profiles induced from $g$ do not form a convergent sequence. As the main result of this paper, we reestablish this convergence result by focusing on a class of large games that satisfy a continuity condition. In particular, Theorem 1 shows that for any large game with a convergent sequence of finite-player games, under this continuity condition, the sequence of obvious strategy profiles forms a convergent sequence of approximate symmetric equilibria. Moreover, such a continuity condition is satisfied in most large games discussed in the literature, including large games with finite characteristics, and continuous large games with continuous equilibria.
Since an obvious strategy profile typically involves randomized strategies, Theorem 1 only guarantees the ex ante asymptotic optimality of such profiles. Therefore, it is natural to consider the ex post equilibrium property of the obvious strategy profile after the resolution of uncertainty. As the second main result of this paper, Theorem 2 shows that, under the same continuity condition, the sequence of realized obvious strategy profiles forms a convergent sequence of approximate equilibria, with probability approaching one. Therefore, based on the concept of obvious strategy profile and the main theorems, we have provided a complete answer to our initial question: players in a large finite game benefit from the idealization process by playing the obvious strategy profile.
Our main results provide a theoretical explanation for human behavior in large strategic interactions. For example, people increasingly use traffic apps, such as Google Maps, to select routes before traveling. In most traffic apps, a driver must input an origin point, a destination point, and the type of vehicle they drive (e.g., car, bus, or truck). The app then calculates the expected travel times for different feasible paths based on idealized models and historical traffic data. Finally, it recommends several optimal choices to the driver based on its calculations and the vehicle type. Typically, a driver will randomly select among the optimal choices recommended by the app, which can be considered an obvious strategy profile. As there are many drivers in the traffic network, each driver’s choice becomes asymptotically optimal not only ex ante but also ex post, with high probability. [^5] There are many other related examples beyond path selection, such as buyers’ behavior in a large online marketplace. For additional examples of symmetric equilibrium in large economies, see Tirole (1988) and Spiegler (2006, 2016).
Then we introduce some results related to Theorem 1. Proposition 1 demonstrates that for a sequence of finite-player games converging to a large game $G$, every Nash equilibrium of $G$ can be approximated by a sequence of approximate symmetric equilibria from the corresponding finite games. Moreover, this result cannot be strengthened by replacing approximate equilibria with exact equilibria. We also examine obvious strategy profiles in pure strategies; Corollary 2 shows that under specific conditions, these obvious strategy profiles exist in pure strategies and form a convergent sequence.
The remainder of this paper is organized as follows. In Section 2, we introduce models of large games and finite-player games, along with the concepts of Nash equilibrium and approximate equilibrium. Our main results are presented in Section 3. Section 4 discuss some related results and reviews related literature. Finally, all the proofs are collected in Section 5.
2 The basic model
In this section, we introduce some notations and basic definitions of large games and large finite-player games. In such games, all the players have a common compact set of actions, and each player’s payoff depends on her own choice as well as the action distribution induced by all the players’ choices.
2.1 Large games
A large game is defined as follows. Let $(I,\mathcal{I},\lambda)$ be an atomless probability space denoting the set of players, [^6] Throughout this paper, we follow the convention that a probability space is complete and countably additive. and $A$ be a compact metric space representing a common action space endowed with the Borel $\sigma$-algebra $\mathcal{B}(A)$. [^7] To simplify the analysis, we restrict attention to large games with common action spaces. All the main results in this paper can be generalized to encompass the setting where players have different feasible action sets. See Remark 1 for more details. The set of Borel probability measures on $A$ is denoted by $\mathcal{M}(A)$. Given an action profile of all the players, the action distribution that specifies the portions of players taking some actions in $A$ (also called a societal summary) can be viewed as an element in $\mathcal{M}(A)$. Each player’s payoff is a bounded continuous function on $A\times\mathcal{M}(A)$, which means that the payoff function continuously depends on her own choice and societal summary.
Let $\mathcal{U}_{A}$ be the space of bounded continuous functions on $A\times\mathcal{M}(A)$ endowed with the sup-norm topology and the resulting Borel $\sigma$-algebra. In a general large game, a player’s characteristic comprises her feasible action set and her payoff function $u_{i}$ (an element of $\mathcal{U}_{A}$). Since players have a common action set $A$ in our setting, $\mathcal{U}_{A}$ can be viewed as the characteristic space for all the players.
A large game $G$ is a measurable function from $(I,\mathcal{I},\lambda)$ to $\mathcal{U}_{A}$, which assigns a payoff function to each player. A pure strategy profile $f$ is a measurable function from $(I,\mathcal{I},\lambda)$ to $A$. Let $\lambda f^{-1}$ be the societal summary (also denoted by $s(f)$), which is the societal action distribution induced by $f$. That is, $s(f)(B)$ ($B$ is a subset of $A$) is the portion of players taking actions in $B$.
[…]
We can now present the equilibrium concepts in large games. The formal definition of a randomized strategy Nash equilibrium is as follows.
Definition 1 (Randomized strategy Nash equilibrium).
A randomized strategy profile $g\colon I\rightarrow\mathcal{M}(A)$ is said to be a randomized strategy Nash equilibrium if for $\lambda$-almost all $i\in I$,
[…]
Therefore, a randomized strategy profile $g$ is a randomized strategy Nash equilibrium if it is optimal for almost every player with respect to the societal summary $s(g)$ in terms of expected payoff. Since for a pure strategy profile $f$, its average action distribution $s(f)$ reduces to the societal action distribution $\lambda f^{-1}$. We have the following definition of pure strategy Nash equilibrium.
Definition 2 (Pure strategy Nash equilibrium).
A pure strategy profile $f\colon I\rightarrow A$ is said to be a pure strategy Nash equilibrium if, for $\lambda$-almost all $i\in I,$
$u_{i}\bigl{(}f(i),\lambda f^{-1}\bigr{)}\geq u_{i}(a,\lambda f^{-1})\text{ forall }a\in A.$
In the definitions of randomized and pure strategy Nash equilibrium, we only require that almost all players make optimal choices. However, stronger concepts can be established by adjusting the strategies of players with measure zero. While a randomized strategy Nash equilibrium is guaranteed to exist in large games, a pure strategy Nash equilibrium may not exist; see, for example, Qiao and Yu (2014).
2.2 Large finite-player games
In this subsection, we introduce a class of large finite-player games. Let a finite probability space $(I^{n},\mathcal{I}^{n},\lambda^{n})$ denote the set of players. Here we assume that $|I^{n}|=n$ and $\mathcal{I}^{n}$ consists of all the subsets of $I^{n}$ (i.e., the power set of $I^{n}$). For simplicity, let $A$ be the common action space for all the players in this game.
Moreover, each player’s payoff function depends on her own choice and the probability distribution on $A$ that is induced from the action profile (i.e., the societal summary). Clearly, the set of such action distributions is a subset of $\mathcal{M}(A)$ that is denoted by
$D^{n}=\Bigl{\{}\tau\in\mathcal{M}(A)\;\Big{|}\;\tau=\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{a_{j}}\text{ where }a_{j}\in A\text{ for all }j\in I^{n}\Bigr{\}},$
where $\delta_{a}$ denotes the Dirac probability measure that assigns probability one to $\{a\}$, for all $a\in A$. Player $i$’s payoff function is then given by a bounded continuous function $u_{i}^{n}\colon A\times\mathcal{M}(A)\to\mathbb{R}$, clearly, $u_{i}^{n}\in\mathcal{U}_{A}$. Thus, a large finite-player game $G^{n}$ can be viewed as a mapping from $I^{n}$ to $\mathcal{U}_{A}$ such that $G^{n}(i)=u_{i}^{n}$ for all $i\in I^{n}$.
In this finite-player game, a pure strategy profile $f^{n}$ is a mapping from $(I^{n},\mathcal{I}^{n},\lambda^{n})$ to $A$. Hence, given a pure strategy profile $f^{n}$, the payoff function for player $i$ is
$u_{i}^{n}(f^{n})=u_{i}^{n}\Bigl{(}f^{n}(i),\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{f^{n}(j)}\Bigr{)},$
here we slightly abuse the notation $u_{i}^{n}(f^{n})$ to denote player $i$’s payoff given the strategy profile $f^{n}$.
Similarly, a randomized strategy is a probability distribution $\mu\in\mathcal{M}(A)$. A randomized strategy profile $g^{n}$ is a mapping from $(I^{n},\mathcal{I}^{n},\lambda^{n})$ to $\mathcal{M}(A)$. Thus, given a randomized strategy profile $g^{n}$, player $i$’s (expected) payoff is
$u_{i}^{n}(g^{n})=\int_{A^{n}}u_{i}^{n}\Bigl{(}a_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{a_{j}}\Bigr{)}\mathbin{\mathop{\otimes}\limits_{j\in I^{n}}}g^{n}(j,{\rm{d}}a_{j}),$
where $\mathbin{\mathop{\otimes}\limits_{j\in I^{n}}}g^{n}(j,{\rm{d}}a_{j})$ is the product probability measure on the product space $A^{n}$. The societal summary induced by $g^{n}$ is $s(g^{n})=\int_{I^{n}}g^{n}(i)\mathrm{d}\lambda^{n}(i)$. Moreover, a randomized strategy profile $g^{n}$ is said to be symmetric if for any two players $i$ and $i^{\prime}$, $g^{n}(i)=g^{n}(i^{\prime})$ whenever $G^{n}(i)=G^{n}(i^{\prime})$. Finally, we state the definitions of randomized strategy Nash equilibrium and $\varepsilon$-Nash equilibrium as follows.
Definition 3 (Randomized strategy Nash equilibrium).
A randomized strategy profile $g^{n}\colon I^{n}\rightarrow\mathcal{M}(A)$ is said to be a randomized strategy Nash equilibrium if for all $i\in I^{n}$,
$u_{i}^{n}(g^{n})\geq u_{i}^{n}(\mu,g^{n}_{-i})\text{ for all }\mu\in\mathcal{M}(A),$
where $(\mu,g^{n}_{-i})$ represents the randomized strategy profile such that player $i$ plays the randomized strategy $\mu$, and player $j$ plays the randomized strategy $g^{n}(j)$ for all $j\in I^{n}\backslash\{i\}$.
Definition 4 ($\varepsilon$-Nash equilibrium).
For any $\varepsilon\geq 0$, a randomized strategy profile $g^{n}\colon I^{n}\to\mathcal{M}(A)$ is said to be an $\varepsilon$-Nash equilibrium if there exists a subset of players $I^{n}_{\varepsilon}\subseteq I^{n}$ such that $\lambda^{n}(I^{n}_{\varepsilon})\geq 1-\varepsilon$ and for all $i\in I^{n}_{\varepsilon}$,
$u_{i}^{n}(g^{n})\geq u_{i}^{n}(\mu,g^{n}_{-i})-\varepsilon\text{ for all }\mu\in\mathcal{M}(A).$
Thus, in an $\varepsilon$-Nash equilibrium, most players choose strategies that are within $\varepsilon$ of their best responses, and only a small portion of players (no more than $\varepsilon$) may obtain higher than $\varepsilon$ by deviation. Clearly, a Nash equilibrium is also an $\varepsilon$-Nash equilibrium ($\varepsilon=0$).
Throughout the rest of this paper, a Nash equilibrium always refers to a randomized strategy Nash equilibrium, and an $\varepsilon$-Nash equilibrium is also called an approximate Nash equilibrium.
3 Main results
In this section, we establish the main results.. Let $G^{n}$ be a large finite-player game with $n$ players. The players in this game use a large game $G$ with a continuum of players as an approximation to simplify the equilibrium analysis. We assume that the players’ characteristic information (i.e., payoff functions) from the finite-player game $G^{n}$ is included in the imagined large game $G$; specifically, we have $G^{n}(I^{n})\subset\operatorname{supp}\lambda G^{-1}$. [^9] Since $G$ is a measurable mapping from the player space $(I,\mathcal{I},\lambda)$ to the characteristic space $\mathcal{U}_{A}$, $\lambda G^{-1}$ is the induced measure on $\mathcal{U}_{A}$. The support of $\lambda G^{-1}$, denoted as $\mathrm{supp}\lambda G^{-1}$, is the smallest closed set $B\in\mathcal{B}(\mathcal{U}_{A})$ such that $\lambda G^{-1}(B)=1$. As we will demonstrate, this assumption allows the players of the large finite-player game $G^{n}$ to implement an equilibrium outcome of the large game $G$ directly.
Suppose that $\tau$ is an action distribution induced by some Nash equilibrium $g$ of large game $G$. This equilibrium action distribution $\tau$ is considered a desirable social choice outcome for players in the large finite game $G^{n}$. As discussed in Section 1, each player may have multiple optimal choices when trying to implement the social choice outcome $\tau$. To address this coordination issue, we adopt a process to symmetrize the equilibrium action distribution. Specifically, we can find a symmetric Nash equilibrium $\widetilde{g}$ of the large game $G$ such that the societal summary of $\widetilde{g}$ also equals $\tau$, i.e., $s(\widetilde{g})=\tau$. This symmetrization result is possible due to the existence of an auxiliary mapping defined as follows.
Definition 5 (Auxiliary mapping).
Given an equilibrium action distribution $\tau$, the auxiliary mapping associated with $\tau$ is a mapping $\overline{g}\colon\mathcal{U}_{A}\to\mathcal{M}(A)$ such that the composition mapping $\overline{g}\circ G\colon I\to\mathcal{M}(A)$ (denoted by $\widetilde{g}$) is a Nash equilibrium of $G$, and satisfies $s(\widetilde{g})=\tau$.
[…]
Figure 1: Auxiliary mapping
Notice that $\overline{g}$ assigns a strategy to each characteristic. Thus, in the strategy profile $\widetilde{g}=\overline{g}\circ G$, players with the same characteristic choose the same strategy. This implies that $\widetilde{g}$ constitutes a symmetric Nash equilibrium with the aggregate action distribution $\tau$.
Lemma 1.
For any equilibrium action distribution $\tau$ of a large game $G$, the auxiliary mapping $\overline{g}$ always exists.
The proof of Lemma 1 is provided in Subsection 5.2.101010Furthermore, it can be shown that given any large game and equilibrium action distribution, the auxiliary mapping is unique. This proof serves as the foundation for the symmetrization process. Clearly, playing a symmetric strategy profile $\widetilde{g}$ avoids coordination issues, as players only need to act according to their characteristics. Since the aggregate action distribution, or societal summary, of $\widetilde{g}$ also equals $\tau$, players can simply focus on the symmetric equilibrium $\widetilde{g}$ to implement $\tau$.
Each player $i$ in the large finite-player game $G^{n}$ has an imagined partner $i^{\prime}$ in the large game $G$ that shares the same characteristics, i.e., $G^{n}(i)=G(i^{\prime})$. [^11] Here we apply the assumption introduced at the beginning of this subsection: $G^{n}(I^{n})\subset\operatorname{supp}\lambda G^{-1}$. A straightforward method for players in $G^{n}$ to implement $\tau$ is for each player to adopt the strategy employed by her imagined partner in $G$, represented as $\widetilde{g}(i^{\prime})=\overline{g}\circ G(i^{\prime})=\overline{g}\circ G^{n}(i)$. This resulting strategy profile in $G^{n}$ is referred to as the obvious strategy profile, which is formally defined as follows.
Definition 6 (Obvious strategy profile).
Given a large finite-player game $G^{n}$, a large game $G$ approximating $G^{n}$, and an auxiliary mapping $\overline{g}$, an obvious strategy profile of $G^{n}$ is defined as $g^{n}=\overline{g}\circ G^{n}$.
[…]
Figure 2: Obvious stratgy profile
Notice that the obvious strategy profile $g^{n}$ is symmetric, as players with the same characteristics select the same strategy determined by $\overline{g}$. Each player in this strategy profile $g^{n}$ chooses the strategy that corresponds to their imagined partner in the strategy profile $\widetilde{g}$ of the large game $G$. We now revisit the motivating example in Section 1 to illustrate $\widetilde{g}$, $\overline{g}$, and $g^{n}$. In that example, $\widetilde{g}(i)=\tfrac{1}{2}\delta_{B}+\tfrac{1}{2}\delta_{H}$ for each player $i$ in the large Beach vs. Hotel game. Since players share a common payoff function $u$ in that game, the characteristic space $G(I)$ is the singleton $u$, and thus $\overline{g}(u)=\tfrac{1}{2}\delta_{B}+\tfrac{1}{2}\delta_{H}$. Therefore, for each finite Beach vs. Hotel game $G^{n}$ with $n$ players, the obvious strategy profile is $g^{n}(i)=\tfrac{1}{2}\delta_{B}+\tfrac{1}{2}\delta_{H}$ for each player $i\in I^{n}$.
Our main theorem examines the equilibrium property of the obvious strategy profile $g^{n}$. According to our analysis in Section 1, $g^{n}$ is not generally an exact Nash equilibrium; therefore, it can be viewed as an $\varepsilon_{n}$-Nash equilibrium for some $\varepsilon_{n}\geq 0$. As $n$ varies, we obtain a sequence of symmetric obvious strategy profiles $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ corresponding to the sequence of finite-player games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$. This leads to a natural question: does the sequence $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ converge? More specifically, does $\varepsilon_{n}\to 0$ as the games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ converge to $G$? To address this question, we first introduce the following concept regarding the convergence of games.
Definition 7.
A sequence of finite-player games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ converges to a large game $G$ if the distributions of characteristics of $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ converge to the distribution of characteristics of $G$. Specifically, this means that $\{\lambda^{n}(G^{n})^{-1}\}_{n\in\mathbb{Z}_{+}}$ converges weakly to $\lambda G^{-1}$.
Now we are ready to present the first main result. In Theorem 1 below, we show that if the auxiliary mapping $\overline{g}$ is almost everywhere continuous on the subset $\mathrm{supp}\lambda G^{-1}\subset\mathcal{U}_{A}$, then the sequence of obvious strategy profiles $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ is a convergent sequence of approximate Nash equilibria.
Theorem 1.
Given a sequence of finite game $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ that converges to a large game $G$, let $\overline{g}$ be an auxiliary mapping of $G$. Define $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ as the sequence of obvious strategy profiles induced by $\overline{g}$. If $\overline{g}$ is almost everywhere continuous on $\mathrm{supp}\lambda G^{-1}$, then there exists a sequence of real numbers $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ such that each $g^{n}$ is an $\varepsilon_{n}$-Nash equilibrium, and $\varepsilon_{n}\to 0$ as $n$ approaches infinity.
The proof of Theorem 1, which is collected in Section 5, comprises two main components and presents two technical contributions. In Subsection 5.1, as the first technical contribution, we establish a theory on the convergence of realized society summaries of randomized strategy profiles. This convergence result is crucial in our proof since an obvious strategy profile $g^{n}$ generally involves randomized strategies. In Subsection 5.3, we make the second technical contribution by decomposing the remaining proof into five steps. We then establish a series of auxiliary inequalities pivotal to the proof. Finally, we finish the proof by using auxiliary inequalities together with the continuity condition on $\overline{g}$.
Notice that the continuity requirement of $\overline{g}$ in Theorem 1 is crucial. Below we present an example showing that if the continuity assumption of $\overline{g}$ is not satisfied, then a sequence of obvious strategy profiles induced by $\overline{g}$ may not be a convergent sequence of approximate symmetric equilibria (i.e., $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ does not converge to 00 as $n$ approaches infinity).
Example 1.
Consider a large game $G$ with the player space $(I,\mathcal{I},\lambda)$ being the Lebesgue unit interval. All the players have a common action set $A=\{0,1\}$. For each player $i\in I$, her payoff function is given by
$u_{i}(a,\tau)=i+\bigl{(}a-\tau(1)\bigr{)}^{2}.$
Hence the large game $G$ can be viewed as a mapping $G(i)=u_{i}$, for all $i\in I$. Let $\mathbb{Q}$ be the set of all rational numbers on $\mathbb{R}$, and $g\colon I\rightarrow\mathcal{M}(A)$ a randomized strategy profile of the game $G$ defined as follows:
$g(i)=\left\{\begin{array}[]{rcl}\delta_{1}&&\textrm{if}\ \ i\in\mathbb{Q}\cap[0,1]\\ i\delta_{0}+(1-i)\delta_{1}&&\textrm{if}\ \ i\in[0,1]\backslash\mathbb{Q}.\\ \end{array}\right.$
Since all the players in this game have different payoff functions, the strategy profile $g$ is a symmetric strategy profile. Moreover, the societal summary $s(g)$ of $g$ is $\int_{I}g(i)\operatorname{d\!}\lambda(i)=\tfrac{1}{2}\delta_{0}+\tfrac{1}{2}\delta_{1}$. Thus we have $u_{i}(0,s(g))=u_{i}(1,s(g))$ for each player $i\in I$, which implies that $g$ is a Nash equilibrium of $G$. Therefore, $g$ is a symmetric Nash equilibrium and its symmetrized strategy profile $\widetilde{g}$ coincides with $g$ itself. That is, $g=\widetilde{g}=\overline{g}\circ G$. Clearly, $G\colon I\to\mathcal{U}_{A}$ is everywhere continuous but $g\colon I\to\mathcal{M}(A)$ is nowhere continuous, hence $\overline{g}\colon\mathcal{U}_{A}\to\mathcal{M}(A)$ is also nowhere continuous.
Then we define a sequence of finite-player games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ that converges to $G$. For each game $G^{n}$, the player space is given by $I^{n}=\{\frac{k}{n}\colon k=1,\ldots,n\}$ endowed with a counting measure $\lambda^{n}$. All the players have a common action set $A=\{0,1\}$. For each player $i\in I^{n}$, her payoff function is given by $u^{n}_{i}(a,\tau)=u_{i}(a,\tau)$, for all $a\in A$ and $\tau\in\mathcal{M}(A).$ Thus we have $G^{n}(i)=u_{i}$ for all $i\in I^{n}$.
For each $n\in\mathbb{Z}_{+}$, we know that $g$ induces an obvious strategy profile $g^{n}$ that is given by $g^{n}(i)=\overline{g}(G^{n}(i))=\overline{g}(G(i))=g(i)$, for all $i\in I^{n}$. Since all the elements of $I^{n}$ are rational numbers, we have $g^{n}(i)=\delta_{1}$ for all $i\in I^{n}$. Thus if $g^{n}$ is viewed as an $\varepsilon_{n}$-Nash equilibrium of $G^{n}$, then $\varepsilon_{n}$ must be no less than $(1-\frac{1}{n})^{2}$, which does not converge to 00.
Despite the counterexample, Theorem 1 remains powerful, as it can be applied to a broad range of large games discussed in the literature. For instance, it is applicable to large games with finite characteristics (i.e., $G(I)$ is a finite subset of $\mathcal{U}_{A}$). Clearly, $\overline{g}$ is almost everywhere continuous when $G(I)$ is a finite set. Consequently, as demonstrated in Corollary 1 below, any sequence of obvious strategy profiles in corresponding finite-player games forms a convergent sequence of approximate symmetric equilibria. Therefore, for players in large finite-player games, adopting the strategies played by their imagined partners in the large game is asymptotically optimal.
Corollary 1.
Let $G\colon(I,\mathcal{I},\lambda)\to\mathcal{U}_{A}$ be a large game such that $G(I)$ is a finite subset of $\mathcal{U}_{A}$, and let $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ be a sequence of finite games converging to $G$. Suppose that $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ is a sequence of obvious strategy profiles induced by $\overline{g}$. Then, there exists a sequence of real numbers $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ such that each $g^{n}$ is an $\varepsilon_{n}$-Nash equilibrium, and $\varepsilon_{n}\to 0$ as $n$ approaches infinity.
Remark 1.
Theorem 1 can be generalized to games with action correspondences, where a player’s characteristic comprises a feasible action set and a payoff function. In this context, one can similarly define the concepts of auxiliary mapping and obvious strategy profile. Furthermore, it can be shown that the sequence of induced obvious strategy profiles is convergent if the auxiliary mapping is almost everywhere continuous on $\mathrm{supp}\lambda G^{-1}$.
Remark 2.
In addition to games with finite characteristics, the continuity condition on $\bar{g}$ is also satisfied in many games with infinite characteristics. For example, when $G\colon I\to\mathcal{U}_{A}$ and $\widetilde{g}\colon I\to\mathcal{M}(A)$ are both continuous mappings, [^12] Here $\mathcal{U}_{A}$ is endowed with the sup-norm topology, while $I$ is equipped with the canonical topology induced by the mapping $G\colon I\to\mathcal{U}_{A}$. Additionally, $\mathcal{M}(A)$ is endowed with the weak topology. it can be observed from Figure 1 that $\overline{g}$ is also continuous. For further discussions on continuous Nash equilibria in continuous games, see Kim (1997). [^13] Kim (1997) adopted the argmax topology on the the characteristic space $\mathcal{U}_{A}$, which differs from the sup-norm topology used in this context.
We conclude this section with the second main result, which examines the equilibrium properties of obvious strategy profiles after the resolution of uncertainty. For any randomized strategy profile $g^{n}\colon I^{n}\to\mathcal{M}(A)$ of the finite-player game $G^{n}$, a pure strategy profile $f^{n}\colon I^{n}\to A$ is said to be a realization of $g^{n}$ if $f^{n}(i)\in\operatorname{supp}g^{n}(i)$ for each $i\in I^{n}$. Let $\Theta_{g^{n}}$ represent the set of all possible realizations of $g^{n}$, and let $\mathbb{P}$ be the induced probability measure on $\Theta_{g^{n}}$. [^14] To be precise, for a randomized strategy profile $g^{n}$, there exist a probability space $(\Omega,\Sigma,\mathbb{P})$ and a mapping $x^{n}\colon I^{n}\times\Omega\to A$ such that the distribution induced from $x^{n}_{i}$ is $g^{n}(i)$. Thus, for each $\omega\in\Omega$, $x^{n}(\omega)$ represents a possible realization of $g^{n}$, and $\mathbb{P}$ is the induced probability measure on the space of realizations.
Now, consider a subset of $\Theta_{g^{n}}$, denoted $\Theta_{g^{n}}^{\varepsilon}$, consisting of all realizations of $g^{n}$ that are (pure strategy) $\varepsilon$-Nash equilibria:
[…]
We are now prepared to present the next main result, demonstrating that the realizations of obvious strategy profiles form a convergent sequence of (pure strategy) approximate equilibria with probability approaching one.
Theorem 2.
Given a sequence of finite game $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ that converges to a large game $G$, let $\overline{g}$ be an auxiliary mapping of $G$. Define $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ as the sequence of obvious strategy profiles induced by $\overline{g}$. If $\overline{g}$ is almost everywhere continuous on $\mathrm{supp}\lambda G^{-1}$, then for any $\varepsilon>0$,
$\lim\limits_{n\to\infty}\mathbb{P}(\Theta_{g^{n}}^{\varepsilon})=1.$
Theorem 2 implies that, for players in a large finite game, adopting the obvious strategy profile will be asymptotically optimal not only ex ante but also ex post, with high probability. The ex post equilibrium property in large finite-player games was first studied by Kalai (2004), who examined the approximate ex post stability of equilibria in large Bayesian games. He showed that if the payoff functions satisfy certain equicontinuity conditions, then realizations of an exact equilibrium in a large Bayesian game are approximately stable with high probability. [^15] While Kalai (2004) focused on games with finite action and type spaces, this result was later generalized to games with general action and type spaces in Carmona and Podczeck (2012) and Deb and Kalai (2015). The main difference between Theorem 2 and Kalai (2004) is that Theorem 2 considers a Nash equilibrium of the limiting large game $G$, which may not yield an exact Nash equilibrium in each finite-player game $G^{n}$. Moreover, as demonstrated in Example 1, the continuity assumption on $\overline{g}$ in Theorem 2 is essential.
4 Further results and discussions
This section is organized into three parts. Subsection 4.1 introduces an equilibrium convergence result, demonstrating that any Nash equilibrium of the large game can be approximated by a convergent sequence of approximate symmetric equilibria derived from the corresponding finite-player games. Subsection 4.2 examines obvious strategy profiles in pure strategies, illustrating that, under specific conditions, such strategy profiles exist and form a convergent sequence. Finally, Subsection 4.3 provides a review of relevant literature from prior studies.
4.1 Convergence of approximate symmetric equilibria
Our main results imply that under certain conditions, every equilibrium distribution $\tau$ of the large game induces a sequence of convergent sequence of approximate symmetric equilibria $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ in the corresponding finite-player games. However, whether the sequence of $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ converges to the equilibrium distribution $\tau$ of the large game remains unknown. Generally, given an equilibrium distribution $\tau$ of the large game, does there exist a sequence of symmetric approximate equilibria $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ of finite-player games such that the societal summaries $s(g^{n})$ converge to $\tau$ as $n$ tends to infinity? We provide an affirmative answer to this question in Proposition 1 below.
Proposition 1.
Given a sequence of finite games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ that converges to $G$, let $\tau$ be an equilibrium action distribution of $G$. Then, there exists a sequence of symmetric strategy profiles $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ of $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ such that $g^{n}$ is an $\varepsilon_{n}$-Nash equilibrium, $\varepsilon_{n}\to 0$, and $s(g^{n})\to\tau$.
Unlike Theorem 1 and Corollary 1, Proposition 1 does not require additional continuity condition or finite-characteristic condition. The proof of Proposition 1 is based on a slight modification of the proof of Theorem 1; therefore, we omit the full detailed proof here but introduce the main idea of our proof instead. Firstly, we consider the case where the continuity condition of $\overline{g}$ holds in $G$. In this situation, we employ the obvious strategy profiles $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ to approximate $\tau$. Indeed, Theorem 1 shows that $\varepsilon_{n}\to 0$, and moreover, the continuity of $\overline{g}$ implies that $s(g^{n})\to\tau$ (see Step 3 of the proof of Theorem 1 in Subsection 5.3). Then for general cases where $\overline{g}$ is not continuous, by using Lusin’s Theorem, [^16] Lusin’s theorem says that for any measurable function $f$, on a metric space $X$ endowed with a finite Borel measure $\lambda$, and for any $\varepsilon>0$, there exists a continuous function $g$ on $X$ such that $\lambda\{x\in X\,|\,g(x)\neq f(x)\}<\varepsilon$. $\overline{g}$ can be approximated by a sequence of continuous functions $\{\overline{g}^{n}\}_{n\in\mathbb{Z}_{+}}$. Following the earlier steps and the proof of Theorem 1, $\overline{g}^{n}\circ G$ can be approximated by a sequence of obvious strategy profiles $\{g^{n,m}\}_{m\in\mathbb{Z}_{+}}$. Finally, employing a diagonal argument, we obtain a sequence of convergent symmetric approximate equilibria that converges to $\tau$.
As a potential enhancement of Proposition 1, a natural question arises: can we strengthen this convergence result by showing that any Nash equilibrium distribution of $G$ can be approximated by a sequence of (exact) Nash equilibria of the corresponding finite-player games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$. However, such an exact lower hemicontinuous property does not generally hold, and Example 2 below serves as a counterexample. [^17] A notable feature of this example is that the limit Nash equilibrium of the large game does not take weakly dominated strategies.
Example 2.
Consider a large game $G$ with the player space $I$ being the Lebesgue unit interval. Players have a common action set $A=\{a,b\}$, and a common payoff function $u$ that is given by:
$u(a,\tau)=\tau(a)\bigl{(}0.9-\tau(a)\bigr{)},\quad u(b,\tau)=0.$
Clearly, there is no weakly dominated strategy, and $g(i)\equiv b$ is a Nash equilibrium. Similarly, we obtain a sequence of finite-player games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ by discretizing $I$.
Claim 1.
The Nash equilibrium $g$ of $G$ cannot be approximated by any sequence of Nash equilibria of $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$.
The detailed proof of Claim 1 is in Subsection 5.5. Here, we provide a brief intuition. Suppose $s(g)=(0,1)$ is a limit point of a sequence $\{s(g^{n})\}_{n\in\mathbb{Z}_{+}}$ where $g^{n}$ is a Nash equilibrium of $G^{n}$. Then for sufficiently large $n$, $s(g^{n})$ is close to $s(g)$, implying that most players in $g^{n}$ choose action $b$ with high probability. After resolving uncertainty, most players indeed choose action $b$. Therefore, players who choose action $b$ with high probability would deviate to action $a$ to obtain higher payoffs, leading to a contradiction. The main technical challenge arises due to the involvement of randomized strategies, resulting in some realized payoffs being less than 0. To address this, we establish a one-to-one matching between negative and positive payoffs, ensuring that the sum of all realized payoffs is positive.
Proposition 1 examines the convergence of Nash equilibria in large games, which is related to results in the literature. However, earlier studies primarily focused on a different question: whether a convergent sequence of Nash equilibria approaches a Nash equilibrium in the limit game. This inquiry concerns the upper hemicontinuity of equilibria, explored in earlier works such as Green (1984) and Housman (1988). Recent advancements on this topic include Khan et al. (2013), Qiao and Yu (2014), and He et al. (2017). In contrast, Proposition 1 in our paper investigates a lower hemicontinuous property of equilibria in large games.
4.2 Obvious strategy profiles in pure strategies
An obvious strategy profile induced from an auxiliary mapping typically involves randomized strategies. A potential extension is to explore the existence of such a strategy profile in pure strategies. Notably, if the symmetric equilibrium $\widetilde{g}$ is a pure strategy profile, then the auxiliary mapping $\overline{g}$ is also pure, as is the resulting obvious strategy profile $g^{n}=\overline{g}\circ G^{n}$. We will examine the atomless condition introduced by Mas-Colell (1984), which ensures the existence of a pure strategy symmetric equilibrium in large games.
Definition 8 (Atomless condition).
A large game $G\colon I\to\mathcal{U}_{A}$ is said to satisfy the atomess condition if the induce distribution $\lambda G^{-1}$ on $\mathcal{U}_{A}$ is atomless.
Mas-Colell (1984, Theorem 2) demonstrated that a pure strategy symmetric equilibrium exists in a large game whenever it satisfies the atomless condition and has a finite action set. [^18] Khan and Sun (1995a, b) generalized this existence result to large games with countable action sets by establishing a pure symmetrization theorem, under the atomless condition. Building on this result and Theorem 1, we derive the following corollary. It is important to note that, throughout the remainder of this subsection, we consider games with a finite action set.
Corollary 2.
Given a sequence of finite game $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ that converges to a large game $G$ satisfying the atomless condition, let $\overline{g}$ be an auxiliary mapping of $G$ induced from a pure strategy symmetric equilibrium $\widetilde{g}$. Define $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ as the sequence of obvious strategy profiles induced by $\overline{g}$. If $\overline{g}$ is almost everywhere continuous on $\mathrm{supp}\lambda G^{-1}$, then there exists a sequence of real numbers $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ such that each $g^{n}$ is an $\varepsilon_{n}$-Nash equilibrium in pure strategy, and $\varepsilon_{n}\to 0$ as $n$ approaches infinity.
The continuity condition is essential in Corollary 2, as the atomless condition is satisfied in Example 1; however, the sequence of obvious strategy profiles does not converge. Corollary 2 can also be proved using the main result from Carmona and Podczeck (2022). Specifically, when $\widetilde{g}$ and $\overline{g}$ are pure, the continuity condition on $\overline{g}$, combined with the mapping theorem (Billingsley (1999, Theorem 2.7)) imply that $\lambda^{n}(G^{n},g^{n})^{-1}=\lambda^{n}(G^{n},\overline{g}\circ G^{n})^{-1}$ converges weakly to $\lambda(G,\overline{g}\circ G)^{-1}=\lambda(G,\widetilde{g})^{-1}$. Consequently, their main theorem asserts that $g^{n}=\overline{g}\circ G^{n}$ constitutes an $\varepsilon_{n}$-Nash equilibrium, where $\varepsilon_{n}\to 0$ as $n$ approaches infinity. However, this proof method is not applicable when $\overline{g}$ is not pure, due to the technical challenges associated with the potential realizations of randomized strategy profiles.
Corollary 2 can be extended to encompass exact symmetric equilibria from topological perspective. Recently, Carmona and Podczeck (2020) demonstrated that when the space of large games is endowed with a strong convergence topology [^19] To be precise, the sequence of games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ converges to $G$ in this topology if (i)$\{\lambda(G^{n})^{-1}\}_{n\in\mathbb{Z}_{+}}$ converges weakly to $\lambda G^{-1}$, and (ii) the sequence of supports of distributions also converges. We refer to this as the “strong convergence topology” to distinguish it from the weak convergence topology. and players’ characteristic space is compact, games with a strict equilibrium constitute a generic (i.e., open and dense) subset of the space of large games. Since a strict equilibrium is both pure and symmetric, the main theorem therein implies that for any sequence of finite games converging to a generic large game in the context of strong convergence topology, there exists a symmetric pure strategy equilibrium $\widetilde{g}$ of $G$ that induces a sequence of convergent symmetric equilibria of $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$. Note that the adoption of strong convergence topology is crucial for this generic result and cannot be replaced by the weak convergence topology employed in the literature and in this paper (see Definition 7).
Another extension of Corollary 2 considers large games with more general action spaces. For instance, in a large game $G$ with a countable action set, we can establish, drawing on the main results from Khan and Sun (1995a, b), that the sequence of obvious strategy profiles forms convergent approximate equilibria when the auxiliary mapping is almost everywhere continuous.
4.3 Related literature
Various progress has been made in large game theory since the foundational papers Schmeidler (1973) and Mas-Colell (1984). Recent developments in large game theory include Kalai (2004), Sun (2006), Khan et al. (2013), Qiao and Yu (2014), Yu (2014), He et al. (2017), Khan et al. (2017), Kalai and Shmaya (2018), Carmona and Podczeck (2020), Khan et al. (2020), Sun et al. (2020), Hellwig (2022), Anderson et al. (2022a, b), and Yang (2022). Beyond theoretical advancements, the large game model is also widely used and studied in applied theory literature. Such papers include Acemoglu and Jensen (2015) in macroeconomics, Yannelis (2009) in large economy, Olszewski and Siegel (2016) in large contest games, Che et al. (2019) in large matching, Barelli et al. (2022) in large elections, and Parise and Ozdaglar (2023) in large network games. We will review the directly related papers in this subsection.
The research most closely related to this paper is by Kalai and Shmaya (2018), who studied a model of incomplete-information repeated games with a large but finite number of players. Given the model’s complexity, they introduced the concept of a (symmetric) imagined-continuum equilibrium to simplify the analysis. In this framework, the societal summary—referred to as the “macro-strategy”—is defined as the arithmetic average of players’ strategies, and individual deviations do not impact this societal summary. Their key findings demonstrated the existence of a symmetric Markov equilibrium in the imagined-continuum game and showed that this equilibrium leads to an approximate equilibrium in the finite-player setting, given that a Lipschitz continuity condition on outcome generating function is satisfied. The key distinctions between Kalai and Shmaya (2018) and this paper are as follows.
First, while both Kalai and Shmaya (2018) and this paper focus on symmetric equilibria in the large continuum game, their motivations differ. Kalai and Shmaya (2018) used symmetry primarily to simplify the characterization of equilibrium, whereas we adopt symmetry to address coordination issues arising from the implementation of the large equilibrium outcome.
Second, in Kalai and Shmaya (2018), players possess the same characteristic (i.e., identical payoff function), resulting in the same strategy being adopted in a symmetric equilibrium in both the continuum and finite games. In this context, the continuity condition required in Theorem 1 is trivially satisfied. In contrast, our model allows for different players to have distinct characteristics. As illustrated in Example 1, the continuity condition on the auxiliary function becomes indispensable in our framework.
Third, the proof of the main results in Kalai and Shmaya (2018) technically relies on a finite action set, whereas our proof accommodates a general action space. Moreover, the payoff function in our model is not necessarily Lipschitz continuous.
This paper builds on prior research concerning symmetric equilibrium in large games, including works by Mas-Colell (1984), Khan and Sun (1995a, b), and Sun et al. (2020). Mas-Colell (1984) established the existence of pure strategy symmetric equilibria in large games with finite action sets and atomless characteristic distributions (see Definition 8). Khan and Sun (1995a, b) expanded on this result, addressing large games with countable action sets. More recently, Sun et al. (2020) explored randomized strategy symmetric equilibria and showed the existence of symmetric equilibria in general large games. These foundational existence results in the literature underpin our main contribution. A key distinction in our study is its focus on a finite number of players, whereas existing research predominantly examines large games involving a continuum of players.
5 Appendix
5.1 Technical preparations
Let $A$ be a compact metric space endowed with its Borel $\sigma$-algebra $\mathcal{B}(A)$, and $d_{A}$ denote the metric on $A$. In this subsection, we introduce two equivalent metrics on $\mathcal{M}(A)$. Based on such metrics, we show that the distance between the sequence of realized society summaries and the sequence of society summaries converges to 0 in probability.
  • Let $\rho$ denote the Prohorov metric on $\mathcal{M}(A)$. That is, for all $\tau,\widetilde{\tau}\in\mathcal{M}(A)$, we have
    $\rho(\tau,\widetilde{\tau})=\inf\bigl{\{}\epsilon>0\colon\tau(B)\leqslant\epsilon+\widetilde{\tau}(B^{\epsilon}),\widetilde{\tau}(B)\leqslant\epsilon+\tau(B^{\epsilon})\text{ for all }B\in\mathcal{B}(A)\bigr{\}},$
    where ${B^{\epsilon}}=\bigl{\{}a\in A\colon d_{A}(a,b)<\epsilon{\text{ for some }}b\in B\bigr{\}}.$
  • Let $\beta$ denote the dual-bounded-Lipschitz metric on $\mathcal{M}(A)$. That is, for all $\tau,\widetilde{\tau}\in\mathcal{M}(A)$, we have
    $\beta(\tau,\widetilde{\tau})=\|\tau-\widetilde{\tau}\|_{{BL}}=\sup\Bigl{\{}\bigl{|}\int_{A}h{\mathrm{d}}(\tau-\widetilde{\tau})\bigr{|}\colon\|h\|_{{B}{L}}\leqslant 1\Bigr{\}},$
    where $h$ is bounded continuous on $A$, $\|h\|_{\infty}=\sup\limits_{a\in A}|h(a)|$, $\|h\|_{\mathrm{L}}=\sup\limits_{a\neq b,a,b\in A}\dfrac{|h(a)-h(b)|}{d_{A}(a,b)}$, and $\|h\|_{{BL}}=\|h\|_{\infty}+\|h\|_{{L}}.$
It is known in the literature that $\rho$ and $\beta$ are equivalent metrics; see, for example, Bogachev (2007, Theorem 8.3.2). Given a sequence $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ where each $g^{n}$ is a randomized strategy profile of game $G^{n}$, let $\{x^{n}_{i}\}_{i\in I^{n},n\in\mathbb{Z}_{+}}$ be a sequence of random variables mapping from a probability space $(\Omega,\Sigma,\mathbb{P})$ to $A$ such that
  • (i)
    for each $i\in I^{n}$ and $n\in\mathbb{Z}_{+}$, the distribution induced from $x^{n}_{i}$ is $g^{n}(i)$;
  • (ii)
    for each $n\in\mathbb{Z}_{+}$, the random variables $\{x^{n}_{i}\}_{i\in I^{n}}$ are pairwise independent.
Lemma 2.
Let $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ be a sequence of finite-player games, and $g^{n}$ be a randomized strategy profile of $G^{n}$ for each $n\in\mathbb{Z}_{+}$. For each $\omega\in\Omega$, let $s(x^{n})(\omega)={\sum_{i\in I^{n}}{{\delta_{{x^{n}_{i}(\omega)}}\lambda^{n}(i)}}}$ be a realized societal summary of the randomized strategy profile $g^{n}$, hence $s(x^{n})$ can be viewed as a random variable from $(\Omega,\Sigma,\mathbb{P})$ to $\mathcal{M}(A)$. Then we have
$\beta\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0\text{ and }\rho\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0\text{ in probability,}$,
where $s(g^{n})=\int_{I^{n}}g^{n}(i)\mathrm{d}\lambda^{n}(i)$.
Proof of Lemma 2.
We divide the proof into two steps. In step 1, we show that for any bounded continuous function $h\colon A\to\mathbb{R}$ with $||h||_{BL}\leq 1$, $\int_{A}h{\rm{d}}\bigl{(}{s(x^{n})-s(g^{n})}\bigr{)}\to 0$ in probability. In step 2, we show that $\beta\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0$ in probability. Finally, by the equivalence of $\rho$ and $\beta$, we obtain that $\rho\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0$ in probability.
Step 1. In this step, we prove that for any bounded and continuous function $h\colon A\to\mathbb{R}$ with $\|h\|_{{BL}}\leq 1$, we have
$\int_{A}h{\rm{d}}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}=\sum\limits_{i\in I^{n}}\lambda^{n}(i)\Bigl{(}h\bigl{(}x^{n}_{i}\bigr{)}-\mathbb{E}\bigl{[}h\bigl{(}x^{n}_{i}\bigr{)}\bigr{]}\Bigr{)}\to 0$ (1)
in probability. Fix any $n\in\mathbb{Z}_{+}$, since $\{x^{n}_{i}\}_{i\in I^{n}}$ are pairwise independent and $h$ is a bounded continuous function, we know that $\{h(x^{n}_{i})\}_{i\in I^{n}}$ are also pairwise independent. By the definition of $\|h\|_{BL}\leq 1$, we have $\|h\|_{\infty}\leq 1$ and hence $-1\leq h(x^{n}_{i})\leq 1$, $\text{var}\bigl{(}h(x^{n}_{i})\bigr{)}\leq 1$, for all $i\in I^{n}$. Moreover, by the independence of $\{h(x^{n}_{i})\}_{i\in I^{n}}$, we have
$\text{var}\Bigl{(}\sum\limits_{i\in I^{n}}h(x^{n}_{i})\lambda^{n}(i)\Bigr{)}=\sum\limits_{i\in I^{n}}\bigl{(}\lambda^{n}(i)\bigr{)}^{2}\text{var}\bigl{(}h(x^{n}_{i})\bigr{)}.$
Since $\mathbb{E}\bigl{[}\sum\limits_{i\in I^{n}}\lambda^{n}(i)h(x^{n}_{i})\bigr{]}=\sum\limits_{i\in I^{n}}\lambda^{n}(i)\mathbb{E}\bigl{[}h(x^{n}_{i})\bigr{]}$, for any $\varepsilon>0$, we have
$\mathbb{P}\Bigl{(}\Bigl{|}\sum_{i\in I^{n}}\lambda^{n}(i)\Bigl{(}h\bigl{(}x^{n}_{i}\bigr{)}-\mathbb{E}\bigl{[}h\bigl{(}x^{n}_{i}\bigr{)}\big{]}\Bigr{)}\Bigr{|}\leq\varepsilon\Bigr{)}$
$\geqslant 1-\frac{\sum\limits_{i\in I^{n}}\bigl{(}\lambda^{n}(i)\bigr{)}^{2}\text{var}\Bigl{(}h\bigl{(}x^{n}_{i}\bigr{)}\Bigr{)}}{\varepsilon^{2}}$
$\geqslant 1-\frac{\sum\limits_{i\in I^{n}}\bigl{(}\lambda^{n}(i)\bigr{)}^{2}}{\varepsilon^{2}}$
$\geqslant 1-\frac{\mathop{\sup}\limits_{j\in I^{n}}\lambda^{n}(j)\sum\limits_{i\in I^{n}}{\lambda^{n}(i)}}{\varepsilon^{2}}$
$\geqslant 1-\frac{\mathop{\sup}\limits_{j\in I^{n}}\lambda^{n}(j)}{\varepsilon^{2}},$ (2)
where the first inequality is due to the Chebyshev’s inequality, and the last inequality follows from the fact that $\sum_{i\in I^{n}}{\lambda^{n}(i)}=1$. Combining with $\sup_{j\in I^{n}}\lambda^{n}(j)\to 0$ as $n\to\infty$, we can finish the proof of Formula (1).
Step 2. In this step, we prove that $\beta\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0$ in probability. According to the proof in step 1, we know that for any finite number $m$, and a sequence of bounded continuous functions $\{{p_{l}}\}_{l=1}^{m}$ with $\|p_{l}\|_{BL}\leq 1$ for all $1\leq l\leq m$, we have
$\sum\limits_{i\in I^{n}}\lambda^{n}(i)\Bigl{(}p_{l}(x^{n}_{i})-\mathbb{E}\bigl{[}p_{l}(x^{n}_{i})\bigr{]}\Bigr{)}\rightarrow 0$ (3)
uniformly in probability for $l\in\{1,2,3,\cdots,m\}$.
Let $E=\{h\colon\|h\|_{BL}\leq 1\}$ be a compact space of bounded continuous functions. Given any $\varepsilon>0$, there exists a finite number $m(\varepsilon)$, and a set of functions denoted by $\{h_{l}\}_{l=1}^{m(\varepsilon)}$ such that
  • (i)
    $h_{1},h_{2},...,h_{m(\varepsilon)}\in E$,
  • (ii)
    for any $h\in E$, $\inf\limits_{1\leq l\leq m(\varepsilon)}\sup\limits_{a\in A}\bigl{|}{h(a)-h_{l}(a)}\bigr{|}<\varepsilon$.
For any $h\in E$, we have
$\quad\Bigl{|}\int_{A}h\mathrm{d}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}\Bigr{|}$
$\leq\inf\limits_{1\leq l\leq m(\varepsilon)}\left\{\Bigl{|}\int_{A}h_{l}\mathrm{d}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}\Bigr{|}+\Bigl{|}\int_{A}(h-h_{l})\mathrm{d}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}\Bigr{|}\right\}$
$\leq\sup\limits_{1\leq l\leq m(\varepsilon)}\Bigl{|}\int_{A}h_{l}\mathrm{d}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}\Bigr{|}+\inf\limits_{1\leq l\leq m(\varepsilon)}\Bigl{|}\int_{A}(h-h_{l})\mathrm{d}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}\Bigr{|}$
$\leq\sup\limits_{1\leq l\leq m(\varepsilon)}\Bigl{|}\int_{A}h_{l}\mathrm{d}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}\Bigr{|}+2\varepsilon,$ (4)
where the first inequality follows from the triangle inequality, and the last inequality follows by $\inf_{1\leq l\leq m(\varepsilon)}\sup_{a\in A}|h(a)-h_{l}(a)|<\varepsilon$. Therefore,
$\sup\limits_{h\in E}\Bigl{|}\int_{A}h\mathrm{d}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}\Bigr{|}\leq\sup\limits_{1\leq l\leq m(\varepsilon)}\Bigl{|}{\int_{A}{h_{l}\mathrm{d}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}}}\Bigr{|}+2\varepsilon.$
To finish the proof of $\beta(s(x^{n}),s(g^{n}))\to 0$ in probability, it suffices to show that for any $\eta>0$,
$\lim\limits_{n\rightarrow\infty}\mathbb{P}\Bigl{(}\beta\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\geq\eta\Bigr{)}=\lim\limits_{n\rightarrow\infty}\mathbb{P}\Bigl{(}\sup\limits_{h\in E}\Bigl{|}\int_{A}h\mathrm{d}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}\Bigr{|}\geq\eta\Bigr{)}=0.$
Pick an $\varepsilon$ such that $0<\varepsilon<\frac{\eta}{2}$, then we only need to show that
$\lim\limits_{n\rightarrow\infty}\mathbb{P}\Bigl{(}\sup\limits_{1\leq l\leq m(\varepsilon)}\Bigl{|}\int_{A}h_{l}\mathrm{d}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}\Bigr{|}\geq-2\varepsilon+\eta\Bigr{)}=0.$
Since
$\mathbb{P}\Bigl{(}\sup\limits_{1\leq l\leq m(\varepsilon)}\Bigl{|}\int_{A}h_{l}\mathrm{d}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}\Bigr{|}\geq-2\varepsilon+\eta\Bigr{)}\leq\sum_{l=1}^{m(\varepsilon)}\mathbb{P}\Bigl{(}\Bigl{|}\int_{A}h_{l}\mathrm{d}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}\Bigr{|}\geq-2\varepsilon+\eta\Bigr{)},$
and by using Formula (3), we conclude that $\mathop{\sup}\limits_{h\in E}\Bigl{|}\int_{A}h\text{d}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}\Bigr{|}\to 0$ in probability. ∎
5.2 Proof of Lemma 1
Let $s(G,g)=\int_{I}\delta_{G(i)}\otimes g(i)\operatorname{d\!}\lambda(i)$ be the joint distribution of $G$ and $g$ on the product space $\mathcal{U}_{A}\times A$. Clearly, the marginal distribution $s(G,g)|_{\mathcal{U}_{A}}=\lambda G^{-1}$. Since both $\mathcal{U}_{A}$ and $A$ are Polish spaces, there exists a family of Borel probability measures $\{\mathcal{S}(u,\cdot)\}_{u\in\mathcal{U}_{A}}$ in $\mathcal{M}(A)$, which represents the disintegration of $s(G,g)$ with respect to $\lambda G^{-1}$ on $\mathcal{U}_{A}$. Let $\overline{g}\colon\mathcal{U}_{A}\to\mathcal{M}(A)$ be defined such that $\overline{g}(u)=\mathcal{S}(u)$ for all $u\in\mathcal{U}_{A}$. According to Sun et al. (2020, Lemma 5), the composition mapping $\widetilde{g}=\overline{g}\circ G$ is a Nash equilibrium of $G$ that satisfies $s(\widetilde{g})=s(g)=\tau$. Furthermore, we can ensure that for any player $i\in I$, her strategy $\widetilde{g}(i)$ is a best response with respect to the society summary $s(\widetilde{g})$. This requirement can be met by modifying the strategies of a subset of players with measure zero.
5.3 Proof of Theorem 1
The major difficulty of this proof is to estimate the difference between $u_{i}^{n}(g^{n})$ and $u_{i}^{n}(\mu,g_{-i}^{n})$ for all $\mu\in\mathcal{M}(A)$. We divide this estimation into five steps. Step 1 shows that we can focus on a subset of payoff functions that consists of equicontinuous and uniformly bounded functions. We then estimate the difference between $u_{i}^{n}(\mu,g^{n}_{-i})$ and $u^{n}_{i}\bigl{(}\mu,s(\mu,g^{n}_{-i})\bigr{)}$ for all $\mu\in\mathcal{M}(A)$ in step 2. Such an estimation result enables us to estimate the difference between $u_{i}^{n}(g^{n})$ and $u_{i}^{n}\bigl{(}g^{n}_{i},s(\widetilde{g})\bigr{)}$ in step 3, and the difference between $u_{i}^{n}(\mu,g_{-i}^{n})$ and $u_{i}^{n}\bigl{(}\mu,s(\widetilde{g})\bigr{)}$ for all $\mu\in\mathcal{M}(A)$ in step 4. Finally in step 5, we combine step 4 and step 5 to finish the proof.
Step 1. We show that for any $\varepsilon>0$, there exist a sequence of subsets $S^{n}\subseteq I^{n}$ such that $\lambda^{n}(S^{n})>1-\frac{\varepsilon}{2}$ for all $n\in\mathbb{Z}_{+}$, and $\{u_{i}^{n}\}_{i\in S^{n},n\in\mathbb{Z}_{+}}$ are equicontinuous and uniformly bounded by a constant $M_{\varepsilon}$. For simplicity, let $\mathcal{W}^{n}=\lambda^{n}(G^{n})^{-1}$, and $\mathcal{W}=\lambda G^{-1}$. Since $A\times\mathcal{M}(A)$ is a compact metric space, the space of bounded and continuous functions $\mathcal{U}_{A}$ on $A\times\mathcal{M}(A)$ is a Polish space. By using the Prohorov theorem (Billingsley (1999, Theorem 5.2)), we know that $\{\mathcal{W}^{n}\}_{n\in\mathbb{Z}_{+}}$ is tight, which means that for any $\varepsilon>0$, there exists a compact set $K_{\varepsilon}\subset\mathcal{U}_{A}$ such that $\mathcal{W}^{n}(K_{\varepsilon})>1-\frac{\varepsilon}{2}$ for all $n\in\mathbb{Z}_{+}$.
Since $K_{\varepsilon}$ is a compact set that consists of bounded and continuous functions, the Arzelà-Ascoli theorem (Munkres (2000, Theorem 45.4)) implies that all the functions in $K_{\varepsilon}$ are equicontinuous and uniformly bounded. Let $M_{\varepsilon}$ denote a bound of all the functions in $K_{\varepsilon}$, and $S^{n}=\{i\in I^{n}|u_{i}^{n}\in K_{\varepsilon}\}$ for all $n\in\mathbb{Z}_{+}$. It is clear that $\lambda^{n}(S^{n})>1-\frac{\varepsilon}{2}$.
Step 2. We estimate $\Bigl{|}u_{i}^{n}(\mu,g^{n}_{-i})-u^{n}_{i}\bigl{(}\mu,s(\mu,g^{n}_{-i})\bigr{)}\Big{|}$ in this step. To be precise, we prove that for any $\varepsilon>0$ and any sequence of randomized strategy profiles $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$, there exists $N_{\varepsilon}\in\mathbb{Z}_{+}$ such that for all $n\geq N_{\varepsilon}$, $i\in S^{n}$, $\mu\in\mathcal{M}(A)$, we have
$\left|u_{i}^{n}(\mu,g^{n}_{-i})-u^{n}_{i}\Bigl{(}\mu,\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\Bigr{)}\right|\leq\frac{\varepsilon}{4}.$
Similar to our proof of Lemma 2, we also use the sequence of random variables $\{x^{n}_{i}\}_{i\in I^{n},n\in\mathbb{Z}_{+}}$ to represent all the players’ strategies in $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$. Let $x_{\mu}$ be a random variable that induces the distribution $\mu$, then we have
$u_{i}^{n}(\mu,g^{n}_{-i})=\mathbb{E}\Bigl{[}u_{i}^{n}\Bigl{(}x_{\mu},\lambda^{n}(i)x_{\mu}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)x^{n}_{j}\Bigr{)}\Bigr{]},$
and
$u^{n}_{i}\Bigl{(}\mu,\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\Bigr{)}=\mathbb{E}\Bigl{[}u_{i}^{n}\Bigl{(}x_{\mu},\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\Bigr{)}\Bigr{]}.$
Hereafter, we restrict our attention to functions in $K_{\varepsilon}$. By equicontinuity, we know that for any $\varepsilon>0$, there exists $\eta>0$ such that for any $\tau,\widetilde{\tau}\in\mathcal{M}(A)$ with $\rho(\tau,\widetilde{\tau})\leq\eta$,
$\begin{split}\bigl{|}u(a,\tau)-u(a,\widetilde{\tau})\bigr{|}\leq\frac{\varepsilon}{4(2M_{\varepsilon}+1)}.\end{split}$ (5)
Let $s(x_{\mu},x^{n}_{-i})(\omega)=\lambda^{n}(i)\delta_{x_{\mu}(\omega)}+\sum_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{j}(\omega)}$, and $s(\mu,g^{n}_{-i})=\lambda^{n}(i)\mu+\sum_{j\in I^{n}\backslash\{i\}}\\ \lambda^{n}(j)g^{n}(j)$, for all $\omega\in\Omega$, $\mu\in\mathcal{M}(A)$. The triangle inequality implies that,
$\rho\bigl{(}s(\mu,g^{n}_{-i}),s(x_{\mu},x^{n}_{-i})(\omega)\bigr{)}\leq$ $\rho\bigl{(}s(g^{n}),s(\mu,g^{n}_{-i})\bigr{)}$
$+\rho\bigl{(}s(x^{n})(\omega),s(g^{n})\bigr{)}$
$+\rho\bigl{(}s(x^{n})(\omega),s(x_{\mu},x^{n}_{-i})(\omega)\bigr{)}$ (6)
By the definition of Prohorov metric $\rho$, we know that for any $\omega\in\Omega$, $\mu\in\mathcal{M}(A)$, $i\in I^{n}$,
$\rho\bigl{(}s(x^{n})(\omega),s(x_{\mu},x^{n}_{-i})(\omega)\bigr{)}\leq\sup\limits_{j\in I^{n}}\lambda^{n}(j).$
Since $\sup\limits_{j\in I^{n}}\lambda^{n}(j)\rightarrow 0$, there exists $N_{1}\in\mathbb{Z}_{+}$ such that for any $n\geq N_{1}$, we have $\mathop{\sup}\limits_{j\in I^{n}}\lambda^{n}(j)<\frac{\eta}{4}$. Hence for any $n\geq N_{1}$, $i\in I^{n}$, $\mu\in\mathcal{M}(A)$, $\omega\in\Omega$,
$\rho\bigl{(}s(x^{n})(\omega),s(x_{\mu},x^{n}_{-i})(\omega)\bigr{)}<\frac{\eta}{4}.$ (7)
By the same argument as above, we can see that for any $n\geq N_{1}$, $i\in I^{n}$, $\mu\in\mathcal{M}(A)$,
$\rho\bigl{(}s(g^{n}),s(\mu,g^{n}_{-i})\bigr{)}<\frac{\eta}{4}.$ (8)
Let $\Omega_{1}^{(\frac{\eta}{2},n)}=\bigl{\{}\omega\in\Omega\,|\,\rho\bigl{(}s(x^{n})(\omega),s(g^{n})\bigr{)}<\frac{\eta}{2}\}$ and $\Omega_{2}^{(\frac{\eta}{2},n)}=\Omega\backslash\Omega_{1}^{(\frac{\eta}{2},n)}$. By Lemma 2, for any $\varepsilon>0$, there exists $N_{\varepsilon}\geq N_{1}$ such that for any $n\geq N_{\varepsilon}$,
$\mathbb{P}\Bigl{(}\Omega_{2}^{(\frac{\eta}{2},n)}\Bigr{)}\leq\frac{\varepsilon}{4(2M_{\varepsilon}+1)}.$ (9)
Let
$H_{1}^{(\frac{\eta}{2},n)}=\left|\mathbb{E}\Bigl{[}\Bigl{(}u_{i}^{n}\bigl{(}x_{\mu},s(\mu,g^{n}_{-i})\bigr{)}-u^{n}_{i}\bigl{(}x_{\mu},s(x_{\mu},x^{n}_{-i})\bigr{)}\Bigr{)}\delta_{\Omega_{1}^{(\frac{\eta}{2},n)}}\Bigr{]}\right|,$
and
$H_{2}^{(\frac{\eta}{2},n)}=\left|\mathbb{E}\Bigl{[}\Bigl{(}u_{i}^{n}\bigl{(}x_{\mu},s(\mu,g^{n}_{-i})\bigr{)}-u^{n}_{i}\bigl{(}x_{\mu},s(x_{\mu},x^{n}_{-i})\bigr{)}\Bigr{)}\delta_{\Omega_{2}^{(\frac{\eta}{2},n)}}\Bigr{]}\right|,$
By using the triangle inequality, we have
$\begin{split}\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},s(\mu,g^{n}_{-i})\bigr{)}-u^{n}_{i}\bigl{(}x_{\mu},s(x_{\mu},x^{n}_{-i})\bigr{)}\Bigr{]}\right|\leq H_{1}^{(\frac{\eta}{2},n)}+H_{2}^{(\frac{\eta}{2},n)}.\end{split}$
Then we estimate $H_{1}^{(\frac{\eta}{2},n)}$ and $H_{2}^{(\frac{\eta}{2},n)}$ separately. Note that for any $n\geq N_{\varepsilon}$ and any player $i\in S^{n}$, we have $u_{i}^{n}\in K_{\varepsilon}$.
  • (i)
    By the definition of event $\Omega_{1}^{(\frac{\eta}{2},n)}$ and Inequalities (5), (5.3), (7), and (8), we can see that
    $H_{1}^{(\frac{\eta}{2},n)}\leq\frac{\varepsilon}{4(2M_{\varepsilon}+1)}.$ (10)
  • (ii)
    Since $u_{i}^{n}$ is bounded by $M_{\varepsilon}$, combined with Inequality (9) we have
    $H_{2}^{(\frac{\eta}{2},n)}\leq 2M_{\varepsilon}\frac{\varepsilon}{4(2M_{\varepsilon}+1)}.$ (11)
Combining Inequalities (10) and (11), for any $n\geq N_{\varepsilon}$, we have
$\left|\mathbb{E}\Bigl{[}u^{n}_{i}\bigl{(}x_{\mu},\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u^{n}_{i}\bigl{(}x_{\mu},\lambda^{n}(i)\delta_{x_{\mu}}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}\right|\leq\frac{\varepsilon}{4}.$
That is,
$\left|u_{i}^{n}(\mu,g^{n}_{-i})-u^{n}_{i}\Bigl{(}\mu,\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\Bigr{)}\right|\leq\frac{\varepsilon}{4}.$
Step 3. In this step, we estimate $\Bigl{|}u_{i}^{n}(g^{n})-u_{i}^{n}\bigl{(}g^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{|}$. Similarly, let $\{x^{n}_{i}\}_{i\in I^{n},n\in\mathbb{Z}_{+}}$ represent all the players’ strategies in games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$. The payoff of player $i\in I^{n}$ in game $G^{n}$ with the strategy profile $g^{n}$ can be rewritten as
$u_{i}^{n}(g^{n})=\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}.$
Since there exists some player $i^{\prime}\in I$ such that $G(i^{\prime})=G^{n}(i)$ (i.e., $u_{i^{\prime}}=u^{n}_{i}$), the payoff of player $i^{\prime}$ in large game $G$ with strategy profile $\widetilde{g}$ is equivalent to
$\int_{A}u_{i^{\prime}}\bigl{(}a,s(\widetilde{g})\bigr{)}\widetilde{g}(i^{\prime},\operatorname{d\!}a)=\int_{A}u_{i}^{n}\bigl{(}a,s(\widetilde{g})\bigr{)}g^{n}(i,\operatorname{d\!}a)=\mathbb{E}\bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},s(\widetilde{g})\bigr{)}\bigr{]}.$
By the triangle inequality, we have
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{]}\right|$
$\leq\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}\right|$ (i)
$+\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{]}\right|.$ (ii)
Estimation of (i). By step 2, for any $\varepsilon>0$, there exists a sequence of sets $\{S^{n}\}_{n\in\mathbb{Z}_{+}}$ such that $\lambda^{n}(S^{n})>1-\frac{\varepsilon}{2}$ for all $n\in\mathbb{Z}_{+}$, and a number $N_{\varepsilon}\in\mathbb{Z}_{+}$ such that for all $n\geq N_{\varepsilon}$, $i\in S^{n}$, we have $\Bigl{|}u_{i}^{n}(g^{n})-u_{i}^{n}\bigl{(}g^{n}(i),\sum_{j\in I^{n}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{|}<\frac{\varepsilon}{4}$. That is,
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}\right|<\frac{\varepsilon}{4}.$
Estimation of (ii). Since all the functions in $\{u_{i}^{n}\}_{i\in S^{n},n\in\mathbb{Z}_{+}}$ are uniformly bounded by $M_{\varepsilon}$ and equicontinuous, for the given $\varepsilon>0$, there exists $\eta>0$ such that for any $\tau,\widetilde{\tau}\in\mathcal{M}(A)$ with $\rho(\tau,\widetilde{\tau})<\eta$, we have $|u(a,\tau)-u(a,\widetilde{\tau})|<\frac{\varepsilon}{4(2M_{\varepsilon}+1)}$, for all $a\in A,u\in\{u_{i}^{n}\}_{i\in S^{n},n\in\mathbb{Z}_{+}}$. For any bounded continuous function $h\colon A\to R$, let $\phi(u)=\int_{A}h(a)\overline{g}(u,\operatorname{d\!}a)$ for all $u\in\mathrm{supp}\lambda G^{-1}\subset\mathcal{U}_{A}$. Since $\overline{g}$ is continuous for $\lambda G^{-1}$-almost all $u\in\mathrm{supp}\lambda G^{-1}$, we know that $\phi$ is bounded and continuous for $\lambda G^{-1}$-almost all $u\in\mathrm{supp}\lambda G^{-1}$. Moreover, as $\lambda^{n}(G^{n})^{-1}$ converges weakly to $\lambda G^{-1}$ and $\mathrm{supp}\lambda^{n}(G^{n})^{-1}\subset\mathrm{supp}\lambda G^{-1}$, by Portmanteau Theorem (Klenke (2014, Theorem 13.16)), we have that
$\int_{\mathrm{supp}\lambda G^{-1}}\phi(u)\operatorname{d\!}\lambda^{n}(G^{n})^{-1}(u)\to\int_{\mathrm{supp}\lambda G^{-1}}\phi(u)\operatorname{d\!}\lambda G^{-1}(u).$
By changing of variables, we can see that
$\int_{\mathrm{supp}\lambda G^{-1}}\phi(u)\operatorname{d\!}\lambda^{n}(G^{n})^{-1}(u)=\int_{I^{n}}\phi\bigl{(}G^{n}(i)\bigr{)}\operatorname{d\!}\lambda^{n}(i),$
and
$\int_{\mathrm{supp}\lambda G^{-1}}\phi(u)\operatorname{d\!}\lambda G^{-1}(u)=\int_{I}\phi\bigl{(}G(i)\bigr{)}\operatorname{d\!}\lambda(i).$
According to the definitions of $\overline{g}$, $\widetilde{g}$, and $g^{n}$, we know that
$\phi\bigl{(}G^{n}(i)\bigr{)}=\int_{A}h(a)\overline{g}\bigl{(}G^{n}(i),\operatorname{d\!}a\bigr{)}=\int_{A}h(a)g^{n}(i,\operatorname{d\!}a),$
and
$\phi\bigl{(}G(i)\bigr{)}=\int_{A}h(a)\overline{g}\bigl{(}G(i),\operatorname{d\!}a\bigr{)}=\int_{A}h(a)\widetilde{g}(i,\operatorname{d\!}a).$
Hence for any continuous function $h$, we have
$\int_{I^{n}}\int_{A}h(a)g^{n}(i,\operatorname{d\!}a)\to\int_{I}\int_{A}h(a)\widetilde{g}(i,\operatorname{d\!}a),$
which is equivalent to
$\int_{A}h(a)s(g^{n})(\operatorname{d\!}a)\to\int_{A}h(a)s(\widetilde{g})(\operatorname{d\!}a)$
by changing of variables. Therefore, $s(g^{n})$ converges weakly to $s(\widetilde{g})$, and there exists $\widetilde{N}_{1}\in\mathbb{Z}_{+}$ such that for all $n\geq\widetilde{N}_{1}$, $\rho\bigl{(}s(g^{n}),s(\widetilde{g})\bigr{)}<\eta$. Thus for all $n\geq\widetilde{N}_{1}$, $i\in S^{n}$, we have
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{]}\right|\leq\dfrac{\varepsilon}{4(2M_{\varepsilon}+1)}.$
Combine the estimations of part (i) and part (ii) above, we have that
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\Bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\Bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{]}\right|\leq\frac{\varepsilon}{4}+\dfrac{\varepsilon}{4(2M_{\varepsilon}+1)},$
for all $n\geq\max\{\widetilde{N}_{1},N_{\varepsilon}\}$, $i\in S^{n}$. Therefore, for all $n\geq\max\{\widetilde{N}_{1},N_{\varepsilon}\}$, $i\in S^{n}$, we have
$\Bigl{|}u_{i}^{n}(g^{n})-\int_{A}u_{i}^{n}\bigl{(}a,s(\widetilde{g})\bigr{)}g^{n}(i,\operatorname{d\!}a)\Bigr{|}\leq\frac{\varepsilon}{4}+\dfrac{\varepsilon}{4(2M_{\varepsilon}+1)}.$
That is,
$\Bigl{|}u_{i}^{n}(g^{n})-u_{i}^{n}\bigl{(}g^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{|}\leq\frac{\varepsilon}{4}+\dfrac{\varepsilon}{4(2M_{\varepsilon}+1)}.$
Step 4. We estimate $\Bigl{|}u_{i}^{n}(\mu,g_{-i}^{n})-u_{i}^{n}\bigl{(}\mu,s(\widetilde{g})\bigr{)}\Bigr{|}$ in this step. By using the triangle inequality, we have that
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\delta_{x_{\mu}}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{i}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},s(\widetilde{g})\bigr{)}\Bigr{]}\right|$
$\leq\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\delta_{x_{\mu}}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}\right|$ (I)
$+\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\sum\limits_{j\in I^{n}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}\right|$ (II)
$+\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\sum\limits_{j\in I^{n}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},s(\widetilde{g})\bigr{)}\Bigr{]}\right|$ (III)
for all $\mu\in\mathcal{M}(A)$, where $x_{\mu}$ is the random variable which induces the distribution $\mu$. The estimation of part (I) is the same as part (i) in step 3, while the estimation of part (III) is the same as part (ii) in step 3. Hence there exists $\widetilde{N}_{2}\in\mathbb{Z}_{+}$ such that for all $n\geq\widetilde{N}_{2}$, $i\in S^{n}$, $\mu\in\mathcal{M}(A)$,
[…]
Below we estimate part (II). By the definition of the Prohorov metric $\rho$, we have that for any $\mu\in\mathcal{M}(A)$, $i\in I^{n}$,
$\rho\bigl{(}s(g^{n}),s(\mu,g^{n}_{-i})\bigr{)}\leq\sup\limits_{j\in I^{n}}\lambda^{n}(j).$
Since $\sup_{j\in I^{n}}\lambda^{n}(j)\rightarrow 0$, there exists $\widetilde{N}_{3}\in\mathbb{Z}_{+}$ such that for any $n\geq\widetilde{N}_{3}$, $\sup_{j\in I^{n}}\lambda^{n}(j)<\eta$. Thus, for any $n\geq\widetilde{N}_{3}$, $i\in I^{n}$, $\mu\in\mathcal{M}(A)$, $\rho\bigl{(}s(g^{n}),s(\mu,g^{n}_{-i})\bigr{)}<\eta.$ Recall that for all $i\in S^{n}$, $u_{i}^{n}$ is uniformly bounded by $M_{\varepsilon}$ and equicontinuous. As we proved in part (ii), we have that
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\sum\limits_{j\in I^{n}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}\right|\leq\dfrac{\varepsilon}{4(2M_{\varepsilon}+1)},$
for all $n\geq\widetilde{N}_{3}$, $i\in S^{n}$, $\mu\in\mathcal{M}(A)$. Thus for all $n\geq\max\{\widetilde{N}_{2},\widetilde{N}_{3}\}$, $i\in S^{n}$, $\mu\in\mathcal{M}(A)$,
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\delta_{x_{\mu}}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},s(\widetilde{g})\bigr{)}\Bigr{]}\right|\leq\frac{\varepsilon}{4}+\dfrac{\varepsilon}{2(2M_{\varepsilon}+1)}.$
That is,
$\Bigl{|}u_{i}^{n}(\mu,g_{-i}^{n})-u_{i}^{n}\bigl{(}\mu,s(\widetilde{g})\bigr{)}\Bigr{|}\leq\frac{\varepsilon}{4}+\dfrac{\varepsilon}{2(2M_{\varepsilon}+1)}.$
Step 5. For any $n\geq\max\{N_{\varepsilon},\widetilde{N}_{1},\widetilde{N}_{2},\widetilde{N}_{3}\}$, $i\in S^{n}$, $\mu\in\mathcal{M}(A)$, we have
$\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}$
$\geq\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{]}-\frac{\varepsilon}{4}-\frac{\varepsilon}{4(2M_{\varepsilon}+1)}$
$\geq\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},s(\widetilde{g})\bigr{)}\Bigr{]}-\frac{\varepsilon}{4}-\frac{\varepsilon}{4(2M_{\varepsilon}+1)}$
$\geq\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\delta_{x_{\mu}}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\frac{\varepsilon}{2}-\frac{\varepsilon}{2(2M_{\varepsilon}+1)}.$
The first inequality follows from step 3 and the last inequality follows from step 4. The second inequality is due to the fact that $\widetilde{g}$ is a Nash equilibrium of $G$ such that for any player $i\in I$, her strategy $\widetilde{g}(i)$ is a best response with respect to the society summary $s(\widetilde{g})$, and $G^{n}(i)\in G(I)$ for all $i\in I^{n}$, $n\in\mathbb{Z}_{+}$. Hence we have that
$\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}\geq\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\delta_{x_{\mu}}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\varepsilon,$
for all $n\geq\max\{N_{\varepsilon},\widetilde{N}_{1},\widetilde{N}_{2},\widetilde{N}_{3}\}$, $i\in S^{n}$, $\mu\in\mathcal{M}(A)$. Thus we conclude that $g^{n}$ is an $\varepsilon$-Nash equilibrium of $G^{n}$, for all $n\geq\max\{N_{\varepsilon},\widetilde{N}_{1},\widetilde{N}_{2},\widetilde{N}_{3}\}$.
Given any sequence $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ such that $\varepsilon_{n}>0$ for all $n\in\mathbb{Z}_{+}$ and $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ converges to 00, there exists a strictly increasing sequence $\{\overline{N}_{n}\}_{n\in\mathbb{Z}_{+}}$ such that $\overline{N}_{n}\in\mathbb{Z}_{+}$ for all $n\in\mathbb{Z}_{+}$, and $g^{m}$ is a $\varepsilon_{n}$-Nash equilibrium of $G^{m}$ for all $m\geq\overline{N}_{n}$. Let $\varepsilon_{k}=\varepsilon_{n}$ if $\overline{N}_{n}\leq k<\overline{N}_{n+1}$, for all $n,k\in\mathbb{Z}_{+}$. Thus we have that $\{\varepsilon_{k}\}_{k\in\mathbb{Z}_{+}}$ converges to 00, and $g^{k}$ is an $\varepsilon_{k}$-Nash equilibrium of $G^{k}$ for all $k\geq\overline{N}_{1}$, which completes our proof of Theorem 1.
5.4 Proof of Theorem 2
We need to show that, for any $\varepsilon>0$ and any $\alpha>0$, there exists an integer $N$ such that for any $n\geq N$, the following holds:
$\mathbb{P}(\Theta_{g^{n}}^{\varepsilon})\geq 1-\alpha.$
To simplify the analysis, we will adopt the same notation as in the previous proof. By Sep 1 of the previous proof, $K_{\varepsilon}\subset\mathcal{U}_{A}$ consists of payoff functions that are equicontinuous and uniformly bounded and the set $S^{n}=\{i\in I^{n}|u_{i}^{n}\in K_{\varepsilon}\}$ for all $n\in\mathbb{Z}_{+}$ satisfies $\lambda^{n}(S^{n})>1-\frac{\varepsilon}{2}>1-\varepsilon$.
For each $\omega\in\Omega$, $x^{n}(\omega)$ represents a possible realization of $g^{n}$. Thus, to show that $x^{n}(\omega)$ is an approximate equilibrium, we only need to estimate the following difference:
$u_{i}^{n}\bigl{(}x^{n}(\omega)\bigr{)}-u^{n}_{i}\bigl{(}a,x^{n}_{-i}(\omega)\bigr{)}$
$=$ $u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),s(x^{n})(\omega)\bigr{)}-u^{n}_{i}\Bigl{(}a,s\bigl{(}\delta_{a},x^{n}_{-i}(\omega)\bigr{)}\Bigr{)}$
where $i\in S^{n}$, $a\in A$, and $s\bigl{(}\delta_{a},x^{n}_{-i}(\omega)\bigr{)}=\lambda^{n}(i)\delta_{a}+\sum_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{j}(\omega)}$.
Clearly, for $i\in S^{n}$, the difference between $u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),s(x^{n})(\omega)\bigr{)}$ and $u^{n}_{i}\Bigl{(}a,s\bigl{(}\delta_{a},x^{n}_{-i}(\omega)\bigr{)}\Bigr{)}$ can be decomposed into three parts:
$u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),s(x^{n})(\omega)\bigr{)}-u^{n}_{i}\Bigl{(}a,s\bigl{(}\delta_{a},x^{n}_{-i}(\omega)\bigr{)}\Bigr{)}$
$=$ $u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),s(x^{n})(\omega)\bigr{)}-u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),s(g^{n})\bigr{)}$
$+u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),s(g^{n})\bigr{)}-u_{i}^{n}\bigl{(}a,s(g^{n})\bigr{)}$
$+u_{i}^{n}\bigl{(}a,s(g^{n})\bigr{)}-u^{n}_{i}\Bigl{(}a,s\bigl{(}\delta_{a},x^{n}_{-i}(\omega)\bigr{)}\Bigr{)}.$
By Lemma 2 and Step 1 of the previous proof, $\rho\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0$ in probability, which implies that there exists an integer $N_{1}$ and a subset $\Omega^{*}\subset\Omega$ such that $\mathbb{P}(\Omega^{*})\geq 1-\alpha$, and for each $\omega\in\Omega^{*}$, $i\in S^{n}$ and $n\geq N_{1}$, we have
$u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),s(x^{n})(\omega)\bigr{)}-u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),s(g^{n})\bigr{)}\geq-\tfrac{\varepsilon}{4},$
and
$u_{i}^{n}\bigl{(}a,s(g^{n})\bigr{)}-u^{n}_{i}\Bigl{(}a,s\bigl{(}\delta_{a},x^{n}_{-i}(\omega)\bigr{)}\Bigr{)}\geq-\tfrac{\varepsilon}{4}.$
Thus, it suffices to estimate the term $u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),s(g^{n})\bigr{)}-u_{i}^{n}\bigl{(}a,s(g^{n})\bigr{)}$. Similarly, it can be decomposed into the following three parts:
$u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),s(g^{n})\bigr{)}-u_{i}^{n}\bigl{(}a,s(g^{n})\bigr{)}$
$=$ $u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),s(g^{n})\bigr{)}-u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),g^{n}_{-i}\bigr{)}$
$+u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),g^{n}_{-i}\bigr{)}-u^{n}_{i}(a,g^{n}_{-i})$
$+u^{n}_{i}(a,g^{n}_{-i})-u_{i}^{n}\bigl{(}a,s(g^{n})\bigr{)}.$
According to Step 2 of the previous proof, there exists an integer $N_{2}$ such that for for $i\in S^{n}$ and each $n\geq N_{2}$, we have
$u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),s(g^{n})\bigr{)}-u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),g^{n}_{-i}\bigr{)}\geq-\tfrac{\varepsilon}{6}$
and
$u^{n}_{i}(a,g^{n}_{-i})-u_{i}^{n}\bigl{(}a,s(g^{n})\bigr{)}\geq-\tfrac{\varepsilon}{6}.$
By Theorem 1, we know that $g^{n}$ is approximate equilibrium of $G^{n}.$ Using the approximate equilibrium property of $g^{n}$, there exists an integer $N_{3}$ such that
$u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),g^{n}_{-i}\bigr{)}-u^{n}_{i}(a,g^{n}_{-i})\geq-\tfrac{\varepsilon}{6}$
for $n\geq N_{3}$ and $i\in S^{n}$. Therefore, for each $n\geq\max\{N_{2},N_{3}\}$ and $i\in S^{n}$,
$u_{i}^{n}\bigl{(}x^{n}_{i}(\omega),s(g^{n})\bigr{)}-u_{i}^{n}\bigl{(}a,s(g^{n})\bigr{)}\geq-\tfrac{\varepsilon}{2}.$
In conclusion, let $N=\max\{N_{1},N_{2},N_{3}\}$. Then, for each $\omega\in\Omega^{*}$ and $i\in S^{n}$, we have
$u_{i}^{n}\bigl{(}x^{n}(\omega)\bigr{)}-u^{n}_{i}\bigl{(}a,x^{n}_{-i}(\omega)\bigr{)}\geq-\tfrac{\varepsilon}{4}-\tfrac{\varepsilon}{2}-\tfrac{\varepsilon}{4}=-\varepsilon,$
which implies that $x^{n}(\omega)$ is an $\varepsilon$-Nash equilibrium. Therefore,
$\mathbb{P}(\Theta_{g^{n}}^{\varepsilon})\geq\mathbb{P}(\Omega^{*})\geq 1-\alpha.$
5.5 Proof of Claim 1
The lemma below introduces an inequality that plays an essential role in the proof of Claim 1.
Lemma 3.
For any $n\in\mathbb{N}$. Suppose $\{a_{i}\}_{i=1}^{n}$ and $\{b_{i}\}_{i=1}^{n}$ are sequences of real numbers that satisfy (i) $a_{i},b_{i}\geq 0$ for all $i\in\{1,\cdots,n\}$; and (ii) there exists $r>0$ such that $a_{i}\geq rb_{i}$ for all $i\in\{1,\cdots,n\}$. Then for any integer $k>\frac{n}{2}$,
$\sum_{\{i_{1},i_{2},\cdots,i_{k},j_{1},j_{2},\cdots,j_{n-k}\}}a_{i_{1}}a_{i_{2}}\cdots a_{i_{k}}b_{j_{1}}b_{j_{2}}\cdots b_{j_{n-k}}$
$\geq r^{2k-n}\sum_{\{i_{1},i_{2},\cdots,i_{k},j_{1},j_{2},\cdots,j_{n-k}\}}b_{i_{1}}b_{i_{2}}\cdots b_{i_{k}}a_{j_{1}}a_{j_{2}}\cdots a_{j_{n-k}},$ (12)
where $i_{1}<i_{2}<\cdots<i_{k}$, $j_{1}<j_{2}<\cdots<j_{n-k}$, and $\{i_{1},i_{2},\cdots,i_{k},j_{1},j_{2},\cdots,j_{n-k}\}=\{1,2,\cdots,n\}$.
Proof of Lemma 3.
The proof is divided into two cases.
Case 1. There exists some $i^{*}\in\{1,2,\cdots,n\}$ such that $a_{i^{*}}=0$. Since $b_{i^{*}}\geq 0$ and $a_{i^{*}}\geq rb_{i^{*}}$, we know that $b_{i^{*}}=0$. Since $\{i_{1},i_{2},\cdots,i_{k},j_{1},j_{2},\cdots,j_{n-k}\}=\{1,2,\cdots,n\}$, for each term $a_{i_{1}}a_{i_{2}}\cdots a_{i_{k}}b_{j_{1}}b_{j_{2}}\cdots b_{j_{n-k}}$ in the left hand side of (3), it must contain $a_{i^{*}}$ or $b_{i^{*}}$, hence the left hand side of (3) is 00. The right hand of (3) is also 00 due to the same reason.
Case 2. $a_{i}>0$ for all $i\in\{1,2,\cdots,n\}$. Then we can divide both sides of (3) by $a_{1}a_{2}\cdots a_{n}$. Let $x_{i}=\frac{rb_{i}}{a_{i}}$ for each $i\in\{1,2,\cdots,n\}$. Clearly, $0\leq x_{i}\leq 1$, and (3) is equivalent to
$\sum_{\{j_{1}<j_{2}<\cdots<j_{n-k}\}}x_{j_{1}}x_{j_{2}}\cdots x_{j_{n-k}}\geq\sum_{\{i_{1}<i_{2}<\cdots<i_{k}\}}x_{i_{1}}x_{i_{2}}\cdots x_{i_{k}},$ (13)
Since $k>\frac{n}{2}$ and $0\leq x_{i}\leq 1$ for all $i\in\{1,2,\cdots,n\}$, we know that $k>n-k$, and the following inequality holds:
$x_{j_{1}}x_{j_{2}}\cdots x_{j_{n-k}}\sum_{\{i_{1},\cdots,i_{2k-n}\}\subseteq\{1,2,\cdots,n\}\backslash\{j_{1},j_{2},\cdots,j_{n-k}\}}x_{i_{1}}x_{i_{2}}\cdots x_{i_{2k-n}}$
$\leq x_{j_{1}}x_{j_{2}}\cdots x_{j_{n-k}}\sum_{\{i_{1},\cdots,i_{2k-n}\}\subseteq\{1,2,\cdots,n\}\backslash\{j_{1},j_{2},\cdots,j_{n-k}\}}1$
$=C_{k}^{2k-n}x_{j_{1}}x_{j_{2}}\cdots x_{j_{n-k}},$ (14)
where $C_{k}^{2k-n}=\frac{k!}{(2k-n)!(n-k)!}$.
Then we sum up both sides of (5.5) for all possible $\{j_{1},j_{2},\cdots,j_{n-k}\}\subseteq\{1,2,\cdots,n\}$. The left hand side of the summation would be
$C_{k}^{2k-n}\sum_{\{i_{1}<i_{2}<\cdots<i_{k}\}}x_{i_{1}}x_{i_{2}}\cdots x_{i_{k}}.$
The reason is that each term is a product of $k$ numbers, and the total number of terms is $C_{n}^{n-k}C_{k}^{2k-n}$. Since there are $C_{n}^{k}$ possibly different terms, the coefficient of each term is $C_{n}^{n-k}C_{k}^{2k-n}/C_{n}^{k}=C_{k}^{2k-n}$. On the other hand, it is clear to see that the right hand side of the summation is
$C_{k}^{2k-n}\sum_{\{j_{1}<j_{2}<\cdots<j_{n-k}\}}x_{j_{1}}x_{j_{2}}\cdots x_{j_{n-k}}.$
Therefore, we have
$C_{k}^{2k-n}\sum_{\{j_{1}<j_{2}<\cdots<j_{n-k}\}}x_{j_{1}}x_{j_{2}}\cdots x_{j_{n-k}}\geq C_{k}^{2k-n}\sum_{\{i_{1}<i_{2}<\cdots<i_{k}\}}x_{i_{1}}x_{i_{2}}\cdots x_{i_{k}},$
which completes our proof. ∎
Now, we are able to prove Claim 1. Let $\tau^{*}$ be the equilibrium distribution induced by $g$, i.e., $\tau^{*}(a)=0$, and $\tau^{*}(b)=1$. For each game $G^{n}$, the player set is given by $I^{n}=\{\frac{k}{n}\colon k=1,\cdots,n\}$. Our proof proceeds by contradiction. Suppose there exists a sequence of randomized strategies $\{g^{n}\}_{n=1}^{\infty}$ such that (i) $g^{n}$ is a Nash equilibrium of $G^{n}$ for each $n\in\mathbb{N}$; and (ii) $s(g^{n})$ converges to $\tau^{*}$. Let $I_{1}^{n}=\{i\in I^{n}|g^{n}(i,a)\geq 0.1\}$, and $I_{2}^{n}=\{i\in I^{n}|g^{n}(i,a)<0.1\}$. Since $\tau^{*}(a)=0$, by Markov’s inequality we know that $\lim\limits_{n\to\infty}\frac{\#|I_{1}^{n}|}{n}=0$, where #$\#|I_{1}^{n}|$ represents the number of players in $I_{1}^{n}$. Thus, there exists $N\in\mathbb{N}$ such that
#$\frac{\#|I_{1}^{n}|}{n}\leq 0.01,\text{ for each }n\geq N.$
Fix a game $G^{n}$ with $n\geq N$. For each player $i_{2}\in I_{2}^{n}$, we know that $g^{n}(i_{2},b)>0.9$. Since $g^{n}$ is an equilibrium of game $G^{n}$, we have $u(b,g^{n}_{-{i_{2}}})\geq u(a,g^{n}_{-{i_{2}}})$. Since $u(b,g_{-i_{2}}^{n})=0$ for any action distribution $\tau$, we have $u(b,g_{-i_{2}}^{n})=0$ and hence $u(a,g^{n}_{-i_{2}})\leq 0$. By the definition of $u$,
$u(a,g^{n}_{-i_{2}})=\int_{\prod\limits_{j\in I^{n}\backslash\{i_{2}\}}A}u\Bigl{(}a,\sum\limits_{j\in I^{n}}\frac{1}{n}\delta_{a_{j}}\Bigr{)}\mathbin{\mathop{\otimes}\limits_{j\in I^{n}\backslash\{i_{2}\}}}g^{n}(j,{\rm{d}}a_{j})$
$=\int_{\prod\limits_{j\in I_{1}^{n}}A}\int_{\prod\limits_{k\in I_{2}^{n}\backslash\{i_{2}\}}A}u\Bigl{(}a,\sum\limits_{j\in I_{1}^{n}}\frac{1}{n}\delta_{a_{j}}+\sum\limits_{k\in I_{2}^{n}\backslash\{i_{2}\}}\frac{1}{n}\delta_{a_{k}}+\frac{1}{n}\delta_{a}\Bigr{)}\mathbin{\mathop{\otimes}\limits_{k\in I_{2}^{n}\backslash\{i_{2}\}}}g^{n}(k,{\rm{d}}a_{k})\mathbin{\mathop{\otimes}\limits_{j\in I_{1}^{n}}}g^{n}(j,{\rm{d}}a_{j}).$
Let $\phi^{n}_{i_{2}}(a)=\int_{\prod\limits_{k\in I_{2}^{n}\backslash\{i_{2}\}}A}u\Bigl{(}a,\sum\limits_{j\in I_{1}^{n}}\frac{1}{n}\delta_{a_{j}}+\sum\limits_{k\in I_{2}^{n}\backslash\{i_{2}\}}\frac{1}{n}\delta_{a_{k}}+\frac{1}{n}\delta_{a}\Bigr{)}\mathbin{\mathop{\otimes}\limits_{k\in I_{2}^{n}\backslash\{i_{2}\}}}g^{n}(k,{\rm{d}}a_{k})$. Recall that $A=\{a,b\}$. Letting $m=\#|I_{2}^{n}|-1$, we have
$\phi^{n}_{i_{2}}(a)=\sum_{k=0}^{m}u\Bigl{(}a,\sum\limits_{j\in I_{1}^{n}}\frac{1}{n}\delta_{a_{j}}+\frac{1}{n}\delta_{a}+\frac{m-k}{n}\delta_{a}+\frac{k}{n}\delta_{b}\Bigr{)}p_{k},$ (15)
where
$p_{k}=\sum_{\{x_{1},\cdots,x_{k},y_{1},\cdots,y_{m-k}\}=I_{2}^{n}\backslash\{i_{2}\}}g_{x_{1}}^{n}(x_{1},\operatorname{d\!}b)\cdots g_{x_{k}}^{n}(x_{k},\operatorname{d\!}b)g_{y_{1}}^{n}(y_{1},\operatorname{d\!}a)\cdots g_{y_{m-k}}^{n}(y_{m-k},\operatorname{d\!}a).$
Since $g_{i}^{n}(i,\operatorname{d\!}b)\geq 9g_{i}^{n}(i,\operatorname{d\!}a)$ for all $i\in I_{2}^{n}$, by Lemma 3, we know that $9^{m-2k}p_{k}\leq p_{m-k},$ for any integer $k<\frac{m}{2}$. Now we deal with negative terms in the right hand side of (15). According to the definition of $u$, it is clear that a negative term must satisfy $\frac{m-k+1}{n}+\sum\limits_{j\in I_{1}^{n}}\frac{1}{n}\delta_{a_{j}}>0.9$. Hence $k<\frac{m}{2}$ and $\frac{m-k+1}{n}>0.89$. Consider a one to one matching that maps every negative term $u(\cdot)p_{k}$ to $u(\cdot)p_{m-k}$, then we have
$u\Bigl{(}a,\sum\limits_{j\in I_{1}^{n}}\frac{1}{n}\delta_{a_{j}}+\frac{1}{n}\delta_{a}+\frac{m-k}{n}\delta_{a}+\frac{k}{n}\delta_{b}\Bigr{)}p_{k}+u\Bigl{(}a,\sum\limits_{j\in I_{1}^{n}}\frac{1}{n}\delta_{a_{j}}+\frac{1}{n}\delta_{a}+\frac{k}{n}\delta_{a}+\frac{m-k}{n}\delta_{b}\Bigr{)}p_{m-k}$
$\geq(-0.1)p_{k}+\frac{1}{n}\Bigl{(}0.9-\frac{1}{n}\Bigr{)}p_{m-k}\geq(-0.1)p_{k}+\frac{1}{n}\Bigl{(}0.9-\frac{1}{n}\Bigr{)}9^{m-2k}p_{k}$
$=\Bigl{[}\frac{1}{n}\Bigl{(}0.9-\frac{1}{n}\Bigr{)}9^{m-2k}-0.1\Bigr{]}p_{k}>\Bigl{[}\frac{1}{n}\Bigl{(}0.9-\frac{1}{n}\Bigr{)}9^{2m-2k+2-2-n}-0.1\Bigr{]}p_{k}$
$\geq\Bigl{[}\frac{1}{n}\Bigl{(}0.9-\frac{1}{n}\Bigr{)}9^{1.78n-2-n}-0.1\Bigr{]}p_{k}=\Bigl{[}\frac{1}{n}\Bigl{(}0.9-\frac{1}{n}\Bigr{)}9^{0.78n-2}-0.1\Bigr{]}p_{k}>0,$
for sufficiently large $n$. The first inequality is due to the fact that $u(a,\tau)$ is an increasing function of $\tau(a)$ when $\tau(a)<0.45$, and decreasing when $\tau(a)>0.45$. Thus, the minimum positive payoff when $\tfrac{1}{n}\leq\tau(a)<0.45$ is attained at $\tau(a)=\frac{1}{n}$ with the payoff $\frac{1}{n}(0.9-\frac{1}{n})$, and the global minimum payoff is $-0.1$ when $\tau(a)=1$. The second inequality follows by $p_{m-k}\geq 9^{m-2k}p_{k}$ for any integer $m-k>\frac{m}{2}$. The third and forth inequalities follow from $\frac{m-k+1}{n}>0.89$ and $m<n$.
Hence, every negative term in $\phi_{i_{2}}^{n}(a)$ can be matched with a positive term such that their summation is positive. Hence $\phi_{i_{2}}^{n}(a)>0$ when $n$ is sufficiently large, which implies that $u(a,g_{-i_{2}}^{n})>0$ and contradicts with our assumption $u(a,g_{-i_{2}}^{n})\leq 0$.